### v4

# Identification by non-Gaussianity in structural threshold and smooth transition vector autoregressive models

Identification by non-Gaussianity in structural threshold and smooth transition vector autoregressive models
Savi Virolainen
University of Helsinki
Linear structural vector autoregressive models can be identified statistically without imposing restrictions on the model if the shocks are mutually independent and at most one of them is Gaussian. We show that this result extends to structural threshold and smooth transition vector autoregressive models incorporating a time-varying impact matrix defined as a weighted sum of the impact matrices of the regimes. We also discuss the problem of labelling the shocks, estimation of the parameters, and stationarity the model. The introduced methods are implemented to the accompanying R package sstvars. Our empirical application studies the effects of the climate policy uncertainty shock on the U.S. macroeconomy. In a structural logistic smooth transition vector autoregressive model with two regimes, we find that a positive climate policy uncertainty shock decreases production and increases inflation in times of both low and high economic policy uncertainty, but its inflationary effects are stronger in the periods of high economic policy uncertainty.
Keywords: smooth transition vector autoregression, threshold VAR, nonlinear SVAR, identification by non-Gaussianity, non-Gaussian SVAR
The author thanks Markku Lanne for the helpful discussions and comments as well as the Research Council of Finland for the financial support (Grant 347986).Contact address: Savi Virolainen, Faculty of Social Sciences, University of Helsinki, P. O. Box 17, FI–00014 University of Helsinki, Finland; e-mail: savi.virolainen@helsinki.fi. ORCiD ID: $0$-0002-5075-6821.
1 Introduction
Linear vector autoregressive (VAR) models are a standard tool in empirical macroeconomics due to their ability to effectively capture constant dynamic relationships of the included variables. Macroeconomic systems may, however, exhibit variation in their dynamics, induced by wars, crises, policy shifts, or business cycle fluctuations, for example. Such structural changes cannot be accommodated by linear VAR models, and therefore, nonlinear VAR models that are able capture such dynamics are often employed. Nonlinear structural VAR (SVAR) models are particularly useful, as they facilitate tracing out the causal effects of economic shocks, which may vary in time depending on the initial state of the economy as well as on the sign and size of the shock (Kilian and Lütkepohl, 2017, Chapter 18). To estimate these effects, the economic shocks need to be identified. Conventional identification methods typically employed in nonlinear SVAR models rely on restrictive assumptions, such as zero contemporaneous interactions among some of the variables (see Kilian and Lütkepohl, 2017, Section 18.2.2). Such restrictions with a clear interpretation are advantageous when based on economic reasoning, but often they are economically implausible and imposed just to achieve sufficient identification. To overcome the latter issue, statistical identification methods relying on the statistical properties of the data can be used.
There are two main branches in the statistical identification literature: identification by heteroscedasticity (Rigobon, 2003, Lanne et al., 2010, Bacchiocchi and Fanelli, 2015, Lütkepohl and Netšunajev, 2017, Lewis, 2021, Virolainen, 2025b, and others) and identification by non-Gaussianity (Lanne et al., 2017, Lanne and Luoto, 2021, Lanne et al., 2023, and others). To the best our knowledge, this paper is the first to study identification by non-Gaussianity in nonlinear SVAR models. Under certain statistical conditions, both types of identification methods typically identify the shocks of a linear structural vector autoregressive (SVAR) model without imposing further restrictions. However, identification by heteroscedasticity without additional restrictions (as described in the cited literature) has the major drawback in nonlinear SVAR models that, for each shock, it restricts the relative magnitudes of the impact responses of the variables to stay constant over time (see, e.g., Lütkepohl and Netšunajev, 2017). This is an undesirable property because it is generally preferable to accommodate time-variation in the (relative) impact responses.
This paper contributes to the literature on identification by non-Gaussianity by extending the framework proposed by Lanne et al. (2017) to smooth transition vector autoregressive (STVAR) models (e.g., Anderson and Vahid, 1998), which is major class of nonlinear SVAR models. We also consider the threshold vector autoregressive model of Tsay (1998) as a special case of the STVAR model. The STVAR model can flexibly capture nonlinear data generating dynamics by accommodating multiple regimes and gradual shifts between them, governed by the transition weighs. In contrast to its linear counterpart, the impact matrix of the structural STVAR model should generally allow for time-variation in the impact responses of the variables to the shocks, which complicates identification. Nonetheless, similarly to Lanne et al. (2017), it turns out that identification is achieved when the shocks are mutually independent and at most one of them is Gaussian. We show that under this condition, when the impact matrix of the STVAR model is defined as a weighted sum of the impact matrices of the regimes, the shocks are readily identified up to ordering and signs. The weights of the impact matrices are the transition weights of the regimes, which we assume to be either exogenous, logistic, or of the threshold form, but our results can be extended to other suitable weight functions as well.
In line with the statistical identification literature, external information is required to label the identified structural shocks as economic shocks. Our nonlinear setup has the additional complication that the same shock should be assigned to the same column of the impact matrix across all regimes. Therefore, since our experience shows that the identification can be weak with respect to the ordering and signs of the columns of the regime-specific impact matrices, we recommend adopting a blended identification strategy that combines identification by non-Gaussianity with additional identifying information (cf. Carriero et al., 2024).
Developments related to ours include the new framework of Morioka et al. (2021) called independent innovation analysis (IIA). The IIA estimates the (conditionally) contemporaneously independent innovations of a nonlinear SVAR model without imposing any specific functional form. Under fairly general assumptions, their method consistently estimates the independent innovations up to permutation and scalar component-wise invertible transformations. While the IIA does seem promising, it is not obvious how to apply it in empirical macroeconometrics, as identification only up to such transformations is, in our understanding, insufficient for structural analysis. Moreover, Morioka et al. (2021) assume the innovations follow a distribution in the exponential family, thereby excluding various interesting alternatives. In contrast, our approach facilitates estimation of the parameters of an identified nonlinear SVAR model, enabling structural analysis without restricting the distribution of independent shocks beyond ruling out Gaussian and infinite-variance cases.
Our empirical application studies the macroeconomic effects of climate policy uncertainty (shocks) and considers monthly U.S. data from 1987:4 to 2024:12. Following Khalil and Strobel (2023) and Huang and Punzi (2024), we measure climate policy uncertainty (CPU) with the CPU index (Gavriilidis, 2021), which is constructed based on the amount of newspaper coverage on topics related to CPU. We are interested in studying how the effects of the CPU shock vary depending on the level of economic policy uncertainty (EPU). Therefore, we fit a two-regime structural logistic STVAR model using the first lag of the EPU index (Baker et al., 2016) as the switching variable. We find that a positive CPU shock decreases production and increases inflation in times of both low and high EPU, but its inflationary effects are stronger in the periods of high EPU. Our results are, hence, in line with the previous literature suggesting that a positive CPU shock decreases production (Fried et al., 2022, Khalil and Strobel, 2023, and Huang and Punzi, 2024) and increases inflation (Khalil and Strobel, 2023, and Huang and Punzi, 2024).
The rest of this paper is organized as follows. Section 2 presents the framework of reduced form STVAR models and provides examples of the covered models. Section 3 discusses identification of the shocks in structural STVAR models, presents our results on identification by non-Gaussianity, and discusses the problem of labelling the shocks. Section 4 discusses stationarity of the model and proposes estimating its parameters with a penalized likelihood -based estimator using a three-step procedure. Section 5 presents the empirical application and Section 6 concludes. Appendices provide detailed discussion on ergodic stationarity of the model; proofs for the stated lemmas, propositions, and theorem; details related to the proposed estimation procedure; a Monte Carlo study assessing the performance of the proposed estimator; and details related to the empirical application. Finally, the introduced methods have been implemented to the accompanying R package sstvars (Virolainen, 2025a), which is available via the CRAN repository.
2 Reduced form STVAR models
Let $y_{t}$, $t=1,2,...$, be the $d$-dimensional time series of interest and $\mathcal{F}_{t-1}$ denote the $\sigma$-algebra generated by the random vectors $\{y_{t-j},j>0\}$. We consider STVAR models with $M$ regimes and autoregressive order $p$ assumed to satisfy
$y_{t}$ $=\sum_{m=1}^{M}\alpha_{m,t}\mu_{m,t}+u_{t},\quad u_{t}\sim MD(0,\Omega_{y,t},\nu),$ (2.1)
$\mu_{m,t}$ $=\phi_{m}+\sum_{i=1}^{p}A_{m,i}y_{t-i},\quad m=1,...,M,$ (2.2)
where $\phi_{1},...,\phi_{M}\in\mathbb{R}^{d}$ are the intercept parameters; $A_{1,i},...,A_{M,i}\in\mathbb{R}^{d\times d}$, $i=1,...,p$, are the autoregression matrices; $u_{t}$ is a martingale difference sequence of reduced form innovations; $\Omega_{y,t}$ is the positive definite conditional covariance matrix of $u_{t}$ (conditional on $\mathcal{F}_{t-1}$), which may depend on $\alpha_{1,t},...,\alpha_{M,t}$ and $y_{t-1},...,y_{t-p}$; and $\nu$ collects any other parameters that the distribution of $u_{t}$ depends on. The transition weights $\alpha_{m,t}$ are assumed to be either exogenous (nonrandom) or $\mathcal{F}_{t-1}$-measurable functions of $\{y_{t-j},j=1,...,p\}$, and to satisfy $\sum_{m=1}^{M}\alpha_{m,t}=1$ at all $t$. Thus, the types of transition weights typically used in macroeconomic applications, including exogenous, logistic, and threshold weights discussed below, are accommodated. The transition weights express the proportions of the regimes the process is on at each point of time and determine how the process shifts between them.
It is easy to see that, conditional on $\mathcal{F}_{t-1}$, the conditional mean of the above-described process is $\mu_{y,t}\equiv E[y_{t}|\mathcal{F}_{t-1}]=\sum_{m=1}^{M}\alpha_{m,t}\mu_{m,t}$, a weighted sum the regime-specific means $\mu_{m,t}$ with the weights given by the transition weights $\alpha_{m,t}$. The conditional covariance matrix is $\Omega_{y,t}$, but discussion on its specification is postponed to Section 3. See Hubrich and Teräsvirta (2013) for a survey on STVAR literature, including also specifications more general than ours.
Our framework accommodates various types of transition weights employed in empirical macroeconomics. To provide a few examples, consider first the threshold VAR (TVAR) model (Tsay, 1998), which is obtained from (2.1) and (2.2) by assuming that the transition weights are defined as
$\alpha_{m,t}=\left\{\begin{matrix}1&\text{if}\ \ r_{m-1}<z_{t}\leq r_{m},\\ 0&\text{otherwise},\phantom{aaaaaaa}\end{matrix}\right.$ (2.3)
where $r_{0}\equiv-\infty$, $r_{M}\equiv\infty$, $m=1,...,M$, are threshold parameters, and $z_{t}$ is a switching variable. At each $t$‚ the model defined in Equations (2.1), (2.2) and (2.3) reduces to a linear VAR corresponding to one of the regimes that is determined according to the level of the switching variable $z_{t}$, which can be an exogenous variable, a lagged endogenous variable, or a linear combination of multiple such variables, for example.
Another popular specification is to assume two regimes ($M=2$) and logistic transition weights. In the logistic STVAR model (Anderson and Vahid, 1998), the transition weights vary according to a logistic function as
$\alpha_{1,t}=1-\alpha_{2,t}\ \ \text{and}\ \ \alpha_{2,t}=[1+\exp\{-\gamma(z_{t}-c)\}]^{-1},$ (2.4)
where $z_{t}$ is the switching variable, $c\in\mathbb{R}$ is a location parameter, and $\gamma>0$ is a scale parameter. The location parameter $c$ determines the midpoint of the transition function, i.e., the value of the switching variable when the weights are equal. The scale parameter $\gamma$, in turn, determines the smoothness of the transitions (smaller $\gamma$ implies smoother transitions), and it is assumed strictly positive so that $\alpha_{2,t}$ is increasing in $z_{t}$.
In addition to specifying the transition weights or the error process $u_{t}$ in various ways, different types of models can be obtained by imposing parameter restrictions to the model defined in (2.1)-(2.2). For instance, a model that is linear in the autoregressive (AR) coefficients is obtained by assuming that they are constant across the regimes, i.e., $A_{m,i}=A_{i}$ and $\phi_{m}=\phi$ for all $i=1,...,p$ and $M=1,...,M$. The model then simplifies to
$y_{t}=\phi+\sum_{i=1}^{p}A_{i}y_{t-i}+u_{t},\quad u_{t}\sim MD(0,\Omega_{y,t},\nu).$ (2.5)
The linear model in (2.5) is typically substantially easier to estimate and has much fewer parameters than the fully nonlinear specification in (2.1)-(2.2), making it more feasible for large scale SVAR analyses, but it can still accommodate shifts in the volatility regime through the error process $u_{t}$.
3 Structural STVAR models identified by non-Gaussianity
3.1 Structural STVAR models
A structural STVAR model is obtained from the reduced form model defined in Section 2 by identifying the serially and mutually uncorrelated structural shocks $e_{t}=(e_{1t},...,e_{dt})$ $(d\times 1)$ from the reduced form innovations $u_{t}$. Specifically, the structural shocks are recovered from the reduced form shocks with the transformation
$e_{t}=B_{y,t}^{-1}u_{t},$ (3.1)
where $B_{y,t}$ is an invertible ($d\times d$) impact matrix that governs the contemporaneous relationships of the shocks and may depend on $\alpha_{1,t},...,\alpha_{M,t}$ and $y_{t-1},...,y_{t-p}$. In other words, assuming a unit variance normalization for the structural shocks, the identification problem amounts to finding an impact matrix $B_{y,t}$ such that $\Omega_{y,t}=B_{y,t}B_{y,t}^{\prime}$. The impact matrix is not, however, generally uniquely identified without imposing further restrictions on the model, since there exists many such decompositions of $\Omega_{y,t}$. Various types of restrictions that recover the structural shocks have been proposed in the literature, many of which are discussed in Kilian and Lütkepohl (2017), for example.
Following the literature on statistical identification by non-Gaussianity in the framework proposed by Lanne et al. (2017), our identifying restriction is that the shocks are mutually independent and at most one of them is Gaussian. Under this assumption, unlike in the Gaussian case, different decompositions of the conditional covariance matrix $\Omega_{y,t}=B_{y,t}B_{y,t}^{\prime}$ generally lead to observationally distinct models, effectively identifying a static impact matrix and the shocks (Lanne et al., 2017). We extend this result to a time-varying impact matrix, and therefore, it is useful to parametrize the structural model directly with $B_{y,t}$. Hence, we make the identity $u_{t}=B_{y,t}e_{t}$ explicit in Equation (2.1) as
$y_{t}=\sum_{m=1}^{M}\alpha_{m,t}\left(\phi_{m}+\sum_{i=1}^{p}A_{m,i}y_{t-i}\right)+B_{y,t}e_{t},\quad e_{t}\sim IID(0,I_{d},\nu),$ (3.2)
where $e_{t}$ are independent and identically distributed structural errors with identity covariance matrix and a distribution that may depend on the parameter $\nu$.
To incorporate time-variation in the impact matrix, we specify its functional form. A natural specification for STVAR models is to assume a constant impact matrix for each of the regimes and define the impact matrix of the process to be their weighted sum as
$B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}B_{m},$ (3.3)
where $B_{1},...,B_{M}$ are invertible $(d\times d)$ impact matrices of the regimes. Also the impact matrix $B_{y,t}$ needs to be invertible to ensure positive definiteness of the conditional covariance matrix, which does not automatically follow from the invertibility of $B_{1},...,B_{M}$. Nevertheless, it turns out that $B_{y,t}$ defined in (3.3) is invertible for all $t$ almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, as is stated in the following lemma (which is proven in Appendix B.1).
Lemma 1.
Suppose $B_{1},...,B_{M}$ are invertible $(d\times d)$ matrices and $\alpha_{1,t},...,\alpha_{M,t}$ scalars such that $\sum_{m=1}^{M}\alpha_{m,t}=1$ and $\alpha_{m,t}\geq 0$ for all $t$ and $m=1,...,M$. Then, the matrix $B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ is invertible for all $t$ almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.
The result of Lemma 1 holds almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, which means that the set where the matrices $B_{1},...,B_{M}$ are such that $B_{y,t}$ is singular for some $t$ has a Lebesgue measure zero. This implies that invertible matrices $B_{1},...,B_{M}$ generally lead to an invertible impact matrix $B_{y,t}$, excluding some special cases of $B_{1},...,B_{M}$ for which this result does not hold (e.g., if $B_{2}=-B_{1}$, then $B_{y,t}=0$ when $\alpha_{1,t}=\alpha_{2,t}=0.5$). To further improve intuition, note that the result of Lemma 1 implies that if the matrices $B_{1},...,B_{M}$ are drawn by random from a continuous distribution, the probability of obtaining matrices such that $B_{y,t}$ is singular for some $t$ is zero. In practice, the parameter space should be simply restricted in estimation so that parameter values leading to $B_{y,t}$ that is singular for some $t$ are excluded.
Assuming impact matrix of the form (3.3), the conditional covariance matrix of $y_{t}$, conditional on $\mathcal{F}_{t-1}$, is obtained as
$\Omega_{y,t}=\left(\sum_{m=1}^{M}\alpha_{m,t}B_{m}\right)\left(\sum_{m=1}^{M}\alpha_{m,t}B_{m}\right)^{\prime}=\sum_{m=1}^{M}\alpha_{m,t}^{2}\Omega_{m}+\sum_{m=1}^{M}\sum_{n=1,n\neq m}^{M}\alpha_{m,t}\alpha_{n,t}\Omega_{m,n},$ (3.4)
where $\Omega_{m}\equiv B_{m}B_{m}^{\prime}$ and $\Omega_{m,n}\equiv B_{m}B_{n}^{\prime}$. The conditional covariance matrix of $y_{t}$ can be thus described as a weighted sum of $M^{2}$ matrices with the weights varying in time according to the transition weights $\alpha_{m,t}$, $m=1,...,M$. If the process is completely in one of the regimes, i.e., $\alpha_{m,t}=1$ for some $m\in\{1,...,M\}$, then $\Omega_{y,t}$ reduces to $\Omega_{m}$, implying that this constitutes as the conditional covariance matrix of Regime $m$. When the process is not completely in any of the regimes, $\Omega_{y,t}$ depends on both the regime-specific covariance matrices and the cross terms. Hence, this specification of the structural model is different to the conventional reduced form STVAR specification in which the conditional covariance matrix of $y_{t}$ is a weighted sum of the covariance matrices of the regimes. Nonetheless, our model specification facilitates exploiting non-Gaussianity of the shocks in their identification, as we show in the next section.
3.2 Identification of the shocks by non-Gaussianity
The key identification assumption is that the structural shocks $e_{t}=(e_{1t},...,e_{dt})$ are mutually independent and at most one of them is Gaussian. We also assume that $e_{t}$ is an IID sequence with zero mean and identity covariance matrix as is formally stated in the following assumption.
Assumption 1.
  1. (i)
    The structural error process $e_{t}=(e_{1t},...,e_{dt})$ is a sequence of independent and identically distributed random vectors such that $e_{t}$ is independent of $\mathcal{F}_{t-1}$ and each component $e_{it}$, $i=1,...,d$, follows a continuous distribution with zero mean and unit variance.
  2. (ii)
    The components of $e_{t}=(e_{1t},...,e_{dt})$ are mutually independent and at most one them has a Gaussian marginal distribution.
Note that the assumption of a unit variance is merely a normalization since the variance of the shocks is captured by the impact matrix $B_{y,t}$. Also, due to the time-variation of the impact matrix (3.3), Assumption 1 does not imply that the reduced form innovations $u_{t}$ are independent nor that they are identically distributed.
The next lemma then shows that at each $t$, conditionally on $\mathcal{F}_{t-1}$, the impact matrix $B_{y,t}$ is uniquely identified up to ordering and signs of its columns. This lemma, which is partly similar to Proposition 1 in Lanne et al. (2017), is used in the following sections to show that the impact matrices $B_{1},...,B_{M}$ of the regimes in (3.3) are identified. Moreover, Lemma 2 also shows that if the transition weights are either exogenous and nonrandom, or they follow the threshold (2.3) or the logistic (2.4) process, then also the autoregressive (AR) and weight function parameters are identified. Specific weight functions are assumed here for simplicity, but the result can be extended to other suitable weight functions as well.
Lemma 2.
Consider the STVAR model defined in Equations (3.2) and (3.3) with Assumption 1 satisfied. Then, conditionally on $\mathcal{F}_{t-1}$, at each $t$, $B_{y,t}$ is uniquely identified up to ordering and signs of its columns almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.
Moreover, suppose the observed time series is $y_{-p+1},...,y_{0},y_{1},...,y_{T}$ and that the following conditions hold:
  1. (1)
    the transition weights $\alpha_{m,t}$, $m=1,...,M$, are either exogenous (nonrandom), or they follow the threshold (2.3) or the logistic process (2.4) (with $M=2$ in the latter),
  2. (2)
    $(\phi_{m},\text{vec}(A_{m,1}),...,\text{vec}(A_{m,p}))\neq(\phi_{n},\text{vec}(A_{n,1}),...,\text{vec}(A_{n,p}))$ for all $m\neq n\in\{1,...,M\}$, and
  3. (3)
    there exists $M$ indices $t_{1},...,t_{M}\in\{1,....,T\}$ such that the corresponding coefficient vectors $(\alpha_{1,t_{i}},...,\alpha_{M,t_{i}})$, $i=1,...,M$, are linearly independent.
Then, for models with exogenous weights, the parameters $\phi_{m},A_{m,1},...A_{m,p}$, $m=1,...,M$ are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$. For threshold models (2.3), the parameters $\phi_{m},A_{m,1},...A_{m,p}$, $m=1,...,M$ and $r_{1},...,r_{m-1}$ are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$. For logistic models (2.4), the parameters $\phi_{m},A_{m,1},...A_{m,p}$, $m=1,2$, and $c,\gamma$ are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ and, if the switching variable $z_{t}$ defined in Section 2 is exogenous, $(z_{1},...,z_{T})\in\mathbb{R}^{T}$.
Lemma 2 is proven in Appendix B.2.
We have concluded that at each $t$, conditionally on $\mathcal{F}_{t-1}$, the impact matrix $B_{y,t}$ is unique up to ordering and signs of its columns under Assumption 1. In other words, at each $t$‚ changing the ordering or signs of the columns of $B_{y,t}$ would lead to an observationally equivalent model, but changing $B_{y,t}$ in any other way would lead to an observationally distinct model. The result holds only almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, because $B_{y,t}$ is invertible almost everywhere by Lemma 1. Identification of the AR and weight function parameters, in turn, is achieved only almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ (and $(z_{1},...,z_{T})\in\mathbb{R}^{T}$), because the identification fails in some special cases that comprise a set with Lebesgue measure zero. Condition (3) of Lemma 2 ensures that the transition weights exhibit enough variation to distinguish between different parameter values. If there are only two regimes, this condition is satisfied if transition weights take at least two distinct values.
If the impact matrix is time-invariant as in Lanne et al. (2017), i.e., $B_{y,t}=B$ for some constant matrix $B$, it follows from Lemma 2 that the structural shocks are identified up to ordering and signs. Such statistically identified shocks can then often be labelled as economic shocks based on external information (see Lanne et al., 2017). However, when the impact matrix varies in time, two complications arise. First, as the impact matrix $B_{y,t}$ is not a matrix of constant parameters but a function of parameters, it needs to be shown that the parameters in the specific functional form of $B_{y,t}$ are identified. Second, since the identification of $B_{y,t}$ at each $t$ is only up to ordering and signs of its columns, its unique identification requires that the parameters in the functional form of $B_{y,t}$ are restricted to fix the ordering and signs of the columns of $B_{y,t}$.
To address these complications, consider our functional form of $B_{y,t}$ in Equation (3.3), which defines $B_{y,t}$ as a weighted sum of the impact matrices of the regimes $B_{1},...,B_{M}$ with the weights given by the transition weights $\alpha_{m,t}$, $m=1,...,M$. To fix the ordering and signs of the columns of $B_{y,t}$, such constraints need to be imposed on $B_{1},...,B_{M}$ that under them, the impact matrices $B_{y,t}$ before and after reordering or changing the signs of the columns of any of $B_{1},...,B_{M}$ are observationally distinct at some $t$. However, constraints that are not restrictive in some specifications may be overidentifying in others, while constraints sufficient in some specifications may not be identifying in others, depending on the transition weights and shock distributions. Therefore, it is useful to divide the discussion on identification into the cases of binary (i.e., $\alpha_{m,t}\in\{0,1\}$) and more general transition weights.
3.3 Identification of the shocks in structural TVAR models
Suppose the transition weights are binary, $\alpha_{m,t}\in\{0,1\}$, for all $t$ and $m=1,...,M$, specifically being either exogenous or of the threshold form (2.3). Then at every $t$, the process is completely in one of the regimes, making the impact matrix $B_{y,t}=B_{m}$ for the active regime $m$ with $\alpha_{m,t}=1$. By Lemma 2, $B_{y,t}$ is thereby identified up to ordering and signs of its columns within each regime, and identification requires constraints that fix these orderings and signs for each regime-specific impact matrix $B_{1},...,B_{M}$. For simplicity, and to state the identification result under just-identifying constraints, we assume each shock $e_{it}$, $i=1,...,d$, follows the skewed $t$-distribution of Hansen (1994), which can flexibly capture fat tails and skewness, and nests the Gaussian distribution as a special case. The skewed $t$-distribution is described in Section 4 discussing estimation of the model parameters. [^1] The identification result can be extended to other suitable shock distributions using arguments similar to those in the proof of Proposition 1 in Appendix B.3.
The following proposition, which is proved in Appendix B.3, states the identification result for the TVAR model and models with exogenous binary weights.
Proposition 1.
Consider the STVAR model defined in Equations (3.2) and (3.3) with $B_{1},...,B_{M}$ invertible; Assumption 1 and the conditions of Lemma 2 satisfied; $\alpha_{m,t}$ either of the threshold form (2.3) or exogenous (nonrandom) with $\alpha_{m,t}\in\{0,1\}$ for all $t$ and $m=1,...,M$; and the shocks $e_{it}$, $i=1,...,d$, following the skewed $t$-distribution defined in Equation (4.1) with the parameters $\nu_{i}>2$ and $\lambda_{i}\in(-1,1)$ controlling kurtosis and skewness, respectively. Suppose the observed time series is $y_{-p+1},...,y_{0},y_{1},...,y_{T}$, and assume the following conditions hold:
  1. (a)
    the ordering and signs of the columns of $B_{1}$ are fixed,
  2. (b)
    $(\nu_{i},\lambda_{i})\neq(\nu_{j},\lambda_{j})$ for all $i\neq j\in\{1,...,d\}$,
  3. (c)
    $\lambda_{i}\neq 0$ for all $i=1,...,d$.
Then, the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T})\in\mathbb{R}^{d(T+p)}$.
Condition (a) of Proposition 1 fixes the ordering and signs of $B_{1}$, which effectively fixes the ordering and signs of the columns of $B_{2},...,B_{M}$ under Conditions (b) and (c), respectively. Any fixed ordering and signs of the columns of $B_{1}$ can be chosen without loss of generality (e.g., by making the first nonzero entry in each column is positive and ordering these entries in decreasing magnitude, given that none of them are equal). If Condition (b) fails for some $i,j\in\{1,...,d\}$, the distributions of these shocks are identical, implying that fixing the ordering of the columns of $B_{1}$ does not fix the ordering of the corresponding columns of $B_{2},...,B_{M}$. In such a case, the ordering of the corresponding columns of $B_{2},...,B_{M}$ can be arbitrarily fixed without loss of generality. Similarly, if Condition (c) fails for some $i\in\{1,...,d\}$ (meaning zero skewness), the signs of the corresponding columns of $B_{2},...,B_{M}$ can be arbitrarily fixed without loss of generality. The result of Proposition 1 holds only almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T})\in\mathbb{R}^{d(T+p)}$, because in some special cases that comprise set with a Lebesgue measure zero, changing the ordering of the columns of $B_{2},...,B_{M}$ does not change the joint density of the model.
If Condition (b) or (c) is close to break, identification is weak with respect to the ordering or signs of the corresponding columns of $B_{2},...,B_{M}$, respectively. Somewhat surprisingly, our experience shows that even when Conditions (b) and (c) hold clearly, identification appears to be often weak with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$, at least if the parameters are estimated by the method of maximum likelihood or penalized maximum likelihood (discussed in Section 4). Specifically, there are frequently a number of local maxima of the (penalized) log-likelihood function with (penalized) log-likelihoods close to each other, corresponding to otherwise largely similar parameter estimates but with different the ordering and signs of the columns of $B_{2},...,B_{M}$. As discussed in Section 3.5, this weak identification should be appropriately dealt with when labelling the shocks, for instance, by combining Proposition 1 with additional identifying information.
3.4 Identification of the shocks in structural STVAR models
Having established the identification result for structural TVAR models, we now consider transition weights that exhibit more variation. In particular, the transition weights are here assumed to be either exogenous (nonrandom) or logistic as defined in (2.4), so that the result of Lemma 2 establishing identification of the transition weights can be exploited. Under the conditions of the following proposition (which is proven in Appendix B.4), the matrices $B_{1},...,B_{M}$ are uniquely identified.
Proposition 2.
Consider the STVAR model defined in Equations (3.2) and (3.3) with Assumption 1 and the conditions of Lemma 2 satisfied, and let the transition weights be either exogenous (nonrandom) or follow the logistic process defined in (2.4) (with $M=2$ assumed). Suppose the observed time series is $y_{-p+1},...,y_{0},y_{1},...,y_{T}$, and assume the following conditions hold:
  1. (A)
    the ordering and signs of the columns of $B_{1}$ are fixed, and
  2. (B)
    the vector $(\alpha_{1,t},...,\alpha_{M,t})$ takes a value for some $t$ such that none of its entries is zero.
Then, for models with logistic transition weights, the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ and, if the switching variable $z_{t}$ is exogenous, $(z_{1},...,z_{T})\in\mathbb{R}^{T}$. For models with nonrandom exogenous transition weights, this result holds almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.
Proposition 2 essentially states that if the structural shocks are independent and at most one of them is Gaussian (Assumption 1), the impact matrices $B_{1},...,B_{M}$, and hence, the structural shocks are identified. The result holds only almost everywhere, because the parameters are not identified in some special cases that constitute a set with Lebesgue measure zero (see also Lemma 2).
Condition (A) of Proposition 2 states that the ordering and signs of the columns of $B_{1}$ should be fixed, which fixes the ordering and signs of the columns of $B_{y,t}$. If the shocks follow the skewed $t$-distributions defined in (4.1), any fixed ordering and signs of the columns of $B_{1}$ can be assumed without loss of generality, as reordering the columns would just reorder the shocks and swapping a sign of a column corresponds to swapping the sign of the related shock and skewness parameter. Condition (B) states that the transition weights should, for some $t$, be strictly positive for all of the regimes. If there are only two regimes, the requirement is merely that the regime-switches are not always discrete, in which case Proposition 1 would be relevant. Similarly to the structural TVAR model discussed in Section 3.3, our experience shows that identification appears to be often weak with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$, which should be appropriately dealt with when labelling the shocks, as discussed in the next section.
3.5 Labelling the shocks and blended identification
As in the linear SVAR model of Lanne et al. (2017), the statistically identified structural shocks do not necessarily have an economic interpretation, and labelling them as economic shocks requires external information. Our identification method, nonetheless, facilitates identifying structural shocks with time-varying impact effects without imposing restrictions on the impact responses of the variables. This is a major advantage compared to statistical identification by heteroscedasticity that restricts how the impact responses may vary over time (see Lütkepohl and Netšunajev, 2017).
Labelling the shocks based on the estimates of the impact matrices $B_{1},...,B_{M}$ might not always be straightforward, as the same shock needs to be associated with the same column of the impact matrix $B_{m}$ in all regimes. For example, if a positive supply shock increases output and decreases prices on impact, and it should do so in all regimes, labelling the $i$th shock as the supply shock requires the $i$th column of all $B_{1},...,B_{M}$ to satisfy such signs. However, as discussed in Sections 3.3 and 3.4, identification can be weak with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$, which should be appropriately accounted for when labelling the shocks.
Assuming that the parameters are estimated by the method of maximum likelihood (ML) or penalized maximum likelihood (PML) described in Section 4, weak identification often results in multiple local maxima of the (penalized) log-likelihood function with the related (penalized) log-likelihoods close to each other. If identification is weak with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$, these local solutions often correspond to otherwise largely similar parameter estimates but with different ordering and signs of the columns of $B_{2},...,B_{M}$.
To formally obtain the local solution that facilitates labelling the shock of interest to the same column of the impact matrix in all regimes, we recommend employing a ”blended identification” strategy that combines identification by non-Gaussianity with additional identifying information (cf. Carriero et al., 2024). As a concrete example, if a supply shock should increase output and decrease prices at impact in all regimes, selecting the best local solution where the same column in all $B_{1},...,B_{M}$ satisfies such signs amounts to combining identification by non-Gaussianity with the appropriate short-run sign restrictions. In addition to short-run sign restrictions, also other types of information can be exploited in blended identification, including narrative restrictions and testable zero impact effect restrictions, for instance.
4 Estimation
4.1 Maximum likelihood estimation and ergodic stationarity
The parameters of the structural STVAR models discussed in Section 3 can be estimated by the method of maximum likelihood (ML), which amounts to maximizing the log-likelihood function over the parameter space. To obtain a well-behaving estimator with desirable asymptotic properties such as consistency, the parameter space is often restricted to the region where the model is ergodic stationary. Based on the results of Saikkonen (2008) (see also Kheifets and Saikkonen, 2020), we show in Appendix A (Theorem A.1) that when the transition weights $\alpha_{m,t}$ are $\mathcal{F}_{t-1}$-measurable functions of $\{y_{t-j},j=1,...,p\}$, a sufficient condition for ergodic stationarity is that the joint spectral radius of the companion form AR matrices of the regimes is strictly less than one. But since this condition is computationally demanding to verify (e.g., Chang and Blondel, 2013), it is not very practical to use it to restrict the parameter space in estimation. Therefore, we restrict the parameter space with the more easily verifiable condition that the usual stability condition is satisfied for each of the regimes (Condition A.1 in Appendix A), as it is necessary for the sufficient condition. After estimation, the sufficient condition can be checked for the solutions of interest.
Maximizing the log-likelihood function (presented in Section 4.2) can be challenging in practice due to its high multimodality, induced by the nonlinear dynamics. Imposing the stability condition can make estimation particularly difficult when the data is persistent, as in such cases numerical optimization algorithms frequently gravitate to the boundary of the parameter space, where they perform poorly and are not often even able to converge to a local maximum. To overcome this issue, we propose to allow for unstable estimates and to maximize the penalized log-likelihood function, obtained by adding a penalty term to the log-likelihood function that penalizes parameter values falling outside or close to the boundary of the stability region. In this way, the optimization algorithm can explore the parameter space also outside the stability region, facilitating improved performance, while sufficient penalization eventually steers the optimization algorithm back to the stability region.
Penalized likelihood -based estimation has been previously applied to time series models by Nielsen and Rahbek (2024). They impose nonnegativity constraints in the autoregressive conditional heteroscedasticity model, allowing the estimation algorithm to explore the boundary of the parameter space. Nielsen and Rahbek (2024) use penalization also to determine which parameters should be set to zero.
4.2 The penalized log-likelihood function
Before introducing the penalized log-likelihood function, the log-likelihood function is presented. We assume that each shock $e_{it}$, $i=1,...,d$, follow the skewed $t$-distribution introduced by Hansen (1994) with zero mean, unit variance, $\nu_{i}>2$ degrees of freedom, and skewness controlled by the parameter $\lambda_{i}\in(-1,1)$. The density function of the skewed $t$-distribution, $st(\cdot;\nu_{i},\lambda_{i})$, is given as (Hansen, 1994, Equations (10)-(13)):
$st(e_{it};\nu_{i},\lambda_{i})=b_{i}c_{i}\left(1+\frac{1}{\nu_{i}-2}\left(\frac{b_{i}e_{it}+a_{i}}{1-\mathbbm{1}\{e_{it}<-a_{i}/b_{i}\}\lambda_{i}+\mathbbm{1}\{e_{it}\geq-a_{i}/b_{i}\}\lambda_{i}}\right)^{2}\right)^{-(\nu_{i}+1)/2}$ (4.1)
where $\mathbbm{1}\{e_{it}<-a_{i}/b_{i}\}$ is an indicator function that takes the value one if $e_{it}<-a_{i}/b_{i}$ and zero otherwise, and $\mathbbm{1}\{e_{it}\geq-a_{i}/b_{i}\}$ is an indicator function that takes the value one if $e_{it}\geq-a_{i}/b_{i}$ and zero otherwise. The constants $a_{i},b_{i}$, and $c_{i}$ are defined as $a_{i}=4\lambda_{i}c_{i}\left(\frac{\nu_{i}-2}{\nu_{i}-1}\right)$, $b_{i}=(1+3\lambda_{i}^{2}-a_{i}^{2})^{1/2}$, and $c_{i}=\frac{\Gamma\left(\frac{\nu_{i}+1}{2}\right)}{(\pi(\nu_{i}-2))^{1/2}\Gamma\left(\frac{\nu_{i}}{2}\right)}$, where $\Gamma(\cdot)$ is the Gamma function.
Indexing the observed data as $y_{-p+1},...,y_{0},y_{1},...,y_{T}$, the conditional log-likelihood function, conditional on the initial values $\boldsymbol{y}_{0}=(y_{0},...,y_{-p+1})$, is given as
$L(\boldsymbol{\theta})=\sum_{t=1}^{T}\log\left(|\det(B_{y,t})|^{-1}\prod_{i=1}^{d}st(I_{d,i}^{\prime}B_{y,t}^{-1}(y_{t}-\mu_{y,t});\nu_{i},\lambda_{i})\right),$ (4.2)
where $I_{d,i}$ is the $i$th column of the $d$-dimensional identity matrix, $\mu_{y,t}$ is defined in Section 2 and $B_{y,t}$ in Equation (3.3). The parameters of the model are collected to the vector $\boldsymbol{\theta}=(\phi_{1},...,\phi_{M},$ $\varphi_{1},...,\varphi_{M},\sigma,\alpha,\nu)$, where $\varphi_{m}=(\text{vec}(A_{m,1}),...,\text{vec}(A_{m,p}))$, $m=1,...,M$, $\sigma=(\text{vec}(B_{1}),...,\text{vec}(B_{M}))$, $\alpha$ contains the parameters governing the transition weights, and $\nu=(\nu_{1},...,\nu_{d},\lambda_{1},...,\lambda_{d})$ contains the degree-of-freedom and skewness parameters. [^2] With threshold (2.3) or logistic (2.4) weights, $\alpha=(r_{1},...,r_{M-1})$ or $\alpha=(c,\gamma)$, respectively, whereas with exogenous weights, $\alpha$ is omitted from the parameter vector.
The penalized likelihood function is then defined as
$PL(\boldsymbol{\theta})=L(\boldsymbol{\theta})-P(\boldsymbol{\theta}),$ (4.3)
where $L(\boldsymbol{\theta})$ is defined in (4.2) and $P(\boldsymbol{\theta})\geq 0$ is a penalization term that penalizes the log-likelihood function from parameter values that are close to entering or are in an uninteresting region of the parameter space. To focus on avoiding unstable estimates, we define the penalization term as
$P(\boldsymbol{\theta})=\kappa Td\sum_{m=1}^{M}\sum_{i=1}^{dp}\text{max}\{0,|\rho(\boldsymbol{A}_{m}(\boldsymbol{\theta}))_{i}|-(1-\eta)\}^{2},$ (4.4)
where $|\rho(\boldsymbol{A}_{m}(\boldsymbol{\theta}))_{i}|$ is the modulus of the $i$th eigenvalue of the companion form AR matrix of Regime $m$, the tuning parameter $\eta\in(0,1)$ determines how close to the boundary of the stability region the penalization starts, and $\kappa>0$ determines the strength of the penalization.
Whenever the companion form AR matrix of a regime has eigenvalues greater than $1-\eta$ in modulus, the penalization term (4.4) is greater than zero, and it increases in the modulus of these eigenvalues. Our penalization term incorporates the multiplicative coefficient $Td$ in an attempt to standardize the strength of penalization with respect to the number of observations $T$ and variables $d$, as the magnitude of the log-likelihood typically increases with these quantities. In our empirical application (Section 5), we use the tuning parameter values $\eta=0.05$ and $\kappa=0.2$, thus, penalizing parameter values outside the stability region significantly, while maintaining flexibility near the boundary.
4.3 Three-step estimation procedure
Lanne et al. (2017) propose a three-step estimation procedure for their linear SVAR model identified by non-Gaussianity. In the first step, least squares (LS) estimation produces preliminary estimates of certain parameters, which serve as initial values in numerical optimization algorithms used in the subsequent steps to maximize the log-likelihood function. The nonlinear dynamics of our structural STVAR model, however, add substantial challenges to the estimation problem by inducing a high number of modes to the (penalized) log-likelihood function. To address these challenges, we propose a modified version of the three-step estimation procedure in which multimodality is explicitly taken into account.
Our three-step estimation procedure proceeds with the following steps:
  1. 1.
    Estimate the autoregressive and weight function parameters $\boldsymbol{\theta}_{1}\equiv(\phi_{1},...,\phi_{M},\varphi_{1},...,\varphi_{M},\alpha)$ by the method of (penalized) nonlinear least squares (NLS) (see Appendix C for details). Denote this estimate of $\boldsymbol{\theta}_{1}$ as $\hat{\boldsymbol{\theta}}_{1}^{NLS}$.
  2. 2.
    Estimate the error distribution parameters $\boldsymbol{\theta}_{2}\equiv(\sigma,\nu)$ by (penalized) ML with a genetic algorithm conditionally on the (penalized) NLS estimate $\hat{\boldsymbol{\theta}}_{1}^{NLS}$. Denote the obtained estimate of $\boldsymbol{\theta}_{2}$ as $\hat{\boldsymbol{\theta}}_{2}^{GA}$.
  3. 3.
    Estimate the full parameter vector $\boldsymbol{\theta}=(\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2})$ by (penalized) ML by initializing a gradient based optimization algorithm from $(\hat{\boldsymbol{\theta}}_{1}^{NLS},\hat{\boldsymbol{\theta}}_{2}^{GA})$.
Step 1 obtains initial estimates for the AR and weight function parameters, and it is comparable to Step 1 of Lanne et al. (2017) but with the nonlinearity of the model explicitly accounted for. Notably, the NLS estimation often produces estimates that do not satisfy the usual stability condition for each of the regimes, and thus allowing for instability (but penalizing it) facilitates utilization of the NLS estimates in our three-step procedure. Step 2 is comparable to Step 2 of Lanne et al. (2017), but the multimodality of the (penalized) log-likelihood function is taken into account by making use of a robust estimation algorithm that is able to escape from local maxima. Step 3 finalizes the estimation by initializing a standard gradient based optimization algorithm from the initial estimates obtained from the previous steps, often converging to a nearby local maximum of the (penalized) log-likelihood function.
Due to the high multimodality of the (penalized) log-likelihood function, Steps 2 and 3 should be ran a large number times to improve the reliability of the results. Since multiple estimation rounds are run, the estimation procedure produces a set of estimates, presumably corresponding to different local maximums of the (penalized) log-likelihood function. If there are multiple solutions with (penalized) log-likelihoods almost equal to the greatest found (penalized) log-likelihood, the blended identification procedure discussed in Section 3.5 can be employed to obtain the best local solution that facilitates labelling the shocks of interest. The proposed three-step estimation procedure is implemented in the accompanying R package sstvars (Virolainen, 2025a), which is available via the CRAN repository.
To assess the performance of the penalized maximum likelihood (PML) estimator as well as our three-step estimation procedure, we conduct a small-scale Monte Carlo study, which is discussed in detail in Appendix D. Since estimation is computationally demanding, we consider simple bivariate (structural) TVAR and LSTVAR models with $p=1$ and $M=2$. To evaluate how the PML estimator performs when the AR matrices are in the penalization region, we consider two specifications of both TVAR and LSTVAR models: one where the AR matrices are well inside the stability region and one where they lie clearly in the region where the penalization term is strictly positive.
According to the results (see Tables D.2 and D.3 in Appendix D for details), estimation accuracy is slightly better when the AR matrices are inside the stability region, but even when they are in the penalization region, the estimators of the AR matrices exhibit only minor small-sample bias. Additionally, estimation accuracy appears to be slightly higher for the TVAR models compared to the LSTVAR models. While for most of the parameters there is only very slight bias, we find a notable upward small-sample bias for the scale parameter, for the degrees-of-freedom parameter whose true value is not small, and for the first entry in each impact matrix. The upward biases in the scale and not-small degrees-of-freedom parameters can both be explained by weak discrimination between large and very large estimates, as large values already imply nearly discrete regime-switches and near-Gaussian distributions, respectively. The upward bias in the first entry of each impact matrix, in turn, could be explained by its sensitivity to extreme values induced by the related particularly heavy tailed skewed $t$-distribution. Nonetheless, the bias diminishes for all of the parameters as the sample size increases. Since also the standard deviations of the estimates diminish as the sample size increases, our results are in line with consistency of the PML estimator. Overall, the performance of the proposed PML estimator and the three-step estimation procedure seems reasonable.
5 Empirical application
Our empirical application studies the macroeconomic effects of climate policy uncertainty (CPU). In the related literature, Fried et al. (2022) develop a dynamic general equilibrium model incorporating beliefs about future climate policy, and find that an increased climate policy transition risk shifts investments toward cleaner capital while reducing total investments and output. Khalil and Strobel (2023) reach similar conclusions in their dynamic stochastic general equilibrium (DSGE) model, attributing the effects to financial institutions’ aversion to uncertainty. Huang and Punzi (2024) also use a DSGE model and find that CPU shocks cause firms to delay investments, reducing output and increasing inflation, which is supported by results from their recursively identified Bayesian VAR.
We are interested in studying how the effects of the CPU shock vary depending on the level of general economic policy uncertainty (EPU). Such variation can be accommodated in our structural STVAR model by specifying a logistic transition weight function with the (lagged) level of EPU as the switching variable. This approach allows us to capture potential nonlinearities and regime-dependent effects, providing more detailed results on the macroeconomic effects of the CPU shock than linear SVAR analysis.
5.1 Data
We consider a monthly U.S. dataset consisting of six variables and covering the time period from 1987:4 to 2024:12. Following Khalil and Strobel (2023) and Huang and Punzi (2024), we measure CPU with the climate policy uncertainty index (CPUI) of Gavriilidis (2021), which is constructed based on the amount of newspaper coverage on climate policy uncertainty topics. The availability of the CPUI data determines the beginning of our sample period. In order to control for general economic policy uncertainty, we also include the economic policy uncertainty index (EPUI) (Baker et al., 2016), which is based on newspaper coverage on topics related to economic policy uncertainty as well as on components quantifying the present value of future scheduled tax code expirations and disagreement among professional forecasters over future government purchases and consumer prices.
As a measure of real economic activity, we include the log of industrial production index (IPI), which is detrended by taking first differences. For measuring the price level, we use the logs of consumer price index (CPI) and producer price index (PPI), likewise detrended by taking first differences. Finally, we measure the monetary policy stance with the effective Federal funds rate that is replaced by the Wu and Xia (2016) shadow rate for the zero-lower-bound periods (RATE). The series of the included variables are presented in Figure 1.333The CPUI and EPUI data are retrieved from https://www.policyuncertainty.com, whereas IPI, CPI, PPI, and the Federal funds rate are retrieved from the Federal reserve bank of St. Louis database and the Wu and Xia (2016) shadow rate from the Federal reserve bank of Atlanta’s website.
Figure 1: Monthly U.S. time series covering the period from 1987:4 to 2024:12. From top to bottom, the variables are the climate policy uncertainty index (divided by $10$); economic policy uncertainty index (divided by $10$); the log-differences of industrial production index, consumer price index, and producer price index (multiplied by $100$); and the effective federal funds rate replaced by the Wu and Xia (2016) shadow rate for the zero lower bound periods. The bottom panel shows the transition weights of the High EPU Regime (Regime 2) of the fitted two-regime second-order structural LSTVAR model. The shaded areas indicate the U.S. recessions defined by the NBER.
5.2 The structural LSTVAR model
In order to specify the structural STVAR model (3.2)-(3.3), the transition weights and the distributions of the structural shocks must be specified. We employ the logistic transition weights with two regimes (Anderson and Vahid, 1998) given in Equation (2.4). They facilitate smooth transitions between the regimes and enable the analysis of how the effects of the CPU shock vary depending on the level of EPU, whose first lag is specified as the switching variable. The logistic weight function also has the convenient property that is nests the threshold weight function (Tsay, 1998) as a special case when the scale parameter tends to infinity, effectively incorporating the selection between these two weight functions within estimation.
We estimate the structural LSTVAR model by PML with the three-step procedure proposed in Section 4, assuming that the structural shocks follow the independent skewed $t$-distributions. The skewed $t$-distribution can flexibly capture fat tails and skewness, and as the Gaussian distribution is obtained as a special case, with $\nu_{i}=\infty$ and $\lambda_{i}=0$, the validity of the identifying Assumption 1 can be assessed based on the estimates of these parameters. To avoid overfitting, particularly in the regime accommodating fewer observations, the autoregressive order $p=2$ is selected based on Hannan-Quinn information criterion.
To deal with the problem of weak identification with respect to the ordering and signs of the columns of $B_{2}$, discussed in Section 3.3, 3.4, and 3.5, we implement the following, simplistic blended identification strategy that largely avoids making restrictive assumptions about the effects of the CPU shock. We normalize the first row of $B_{1}$ positive with CPUI ordered first and assume that in both regimes, one of the shocks, deemed as the CPU shock, has greater impact effect on the CPUI than any other shock, and this impact effect has the same sign in both regimes. Then, to support our identification, we check that the recovered CPU shock aligns with several major historical events affecting CPU. In particular, we find that President Trump’s election is preceded by large CPU shocks in October 2016 and 2024 ($17$ and $26$, respectively). The largest negative recovered CPU shock ($-10$), in turn, dates to August 2022, when the Inflation Reduction Act was passed, providing support for the clean energy transition. Among all the shocks recovered from the model, only the one deemed as the CPU shock has signs consistent with the above historical events affecting CPU.
The location parameter estimate is $\hat{c}=14.48$, implying roughly that Regime 1 dominates when the level of EPU is not high, while Regime 2 starts prevailing in times of high EPU. Therefore, we label Regime 1 as Low EPU Regime and Regime 2 as High EPU Regime. The scale parameter estimate is $\hat{\gamma}=16.65$, indicating that transitions between the regimes are somewhat fast (rather than very smooth), which can be observed also from the bottom panel of Figure 1, where the fitted transition weights of High EPU Regime are presented.
The estimates of the degrees-of-freedom and skewness parameters are (2.33,2.10,2.64,3.23,6.54,(2.33,2.10,2.64,3.23,6.54,( 2.33 , 2.10 , 2.64 , 3.23 , 6.54 , 2.68)2.68)2.68 ) and $(0.41,0.17,0.09,0.43,0.25,-0.12)$, respectively. Clearly, none of the shocks is close to Gaussian, suggesting that Assumption 1 is satisfied. Then, we check whether our model satisfies the stationarity condition of Theorem A.1 in Appendix A by computing an upper bound for the joint spectral radius of the companion form AR matrices of the regimes. [^4] We compute the upper bound using the JSR toolbox (Jungers, 2023) in MATLAB. The obtained upper bound is $0.98$, which is strictly less than one, implying that our model satisfies Assumption A.1 (in Appendix A) and is ergodic stationary by Theorem A.1. Based on the residual diagnostics presented in Appendix E, the adequacy of our model to capture the autocorrelation structure, conditional heteroskedasticity, and distribution of the data seems, overall, reasonable.
5.3 Generalized impulse response functions
The effects of the shocks in our structural LSTVAR model may depend on the initial state of the economy as well as on the sign and size of the shock, which makes the conventional way of calculating impulse responses unsuitable. Therefore, we consider the general impulse response function (GIRF) (Koop et al., 1996) that accommodates such features and is defined as
$\text{GIRF}(h,\delta_{i},\mathcal{F}_{t-1})=\text{E}[y_{t+h}|\delta_{i},\mathcal{F}_{t-1}]-\text{E}[y_{t+h}|\mathcal{F}_{t-1}],$ (5.1)
where $h$ is the horizon. The first term on the right side of (5.1) is the expected realization of the process at time $t+h$ conditionally on a structural shock of sign and size $\delta_{i}\in\mathbb{R}$ in the $i$th element of $e_{t}$ at time $t$ and the previous observations. The latter term on the right side is the expected realization of the process conditionally on the previous observations only. The GIRF thus expresses the expected difference in the future outcomes when the $i$th structural shock of sign and size $\delta_{i}$ arrives at time $t$ as opposed to all shocks being random. The GIRFs also facilitate tracking the effects of the shock on the transition weights $\alpha_{m,t}$, $m=1,...,M$, by replacing $y_{t+h}$ with $\alpha_{m,t+h}$ on the right side of Equation (5.1).
Due to the $p$-step Markov property of our structural LSTVAR model, conditioning on the previous $p$ observations, $\boldsymbol{y}_{t-1}=(y_{t-1},...,y_{t-p})$, is equivalent to conditioning on $\mathcal{F}_{t-1}$. To the study the state-dependence of the effects of the CPU shock, we follow a procedure similar to Lanne and Virolainen (2025) and calculate the GIRFs conditional on a given regime dominating when the shock arrives. Specifically, for each Regime $m\in\{1,2\}$, we take all the length $p$ histories $\boldsymbol{y}_{t-1}$ from the data for which the corresponding transition weight $\alpha_{m,t}$ is greater than $0.75$, indicating that Regime $m$ is clearly dominant (there are $366$ such histories for the Low EPU Regime and $80$ for the High EPU Regime). For each such history, we take the corresponding CPU shock recovered from the data, and compute the GIRF using the Monte Carlo algorithm described in Lanne and Virolainen (2025), Appendix B. Finally, GIRFs corresponding to shocks of different sign and size are made comparable by scaling them to correspond to a $5$ point instantaneous increase of CPUI.
Figure 2: Generalized impulse response functions to the identified CPU shock $h=0,1,....,36$ months ahead. From top to bottom, the responses of CPUI, EPUI, IPI, CPI, PPI, the interest rate variable, and transition weights of the High EPU Regime are depicted in each row. The responses of IPI, CPI, and PPI are accumulated to (100×100$\times$100 ×)log-levels. The left column shows the responses to conditionally on the Low EPU Regime dominating when the shock arrives, whereas the right column shows the responses conditionally on the High EPU Regime dominating. All the GIRFs have been scaled so that the instantaneous increase in CPUI is $5$ points. Each GIRF is computed with a Monte Carlo algorithm similar to the one described in Lanne and Virolainen (2025, Appendix B).
The distribution of the GIRFs in each regime is illustrate in a so-called ”shotgun plot” in Figure 2, which depicts each GIRF using a level of opacity such that the darkness of a region displays the concentration of GIRFs in it (cf. Lanne and Virolainen, 2025). The response of CPUI to a positive CPU shock follows a similar pattern in both regimes in the vast majority of the GIRFs, while EPUI temporarily increases in the Low EPU Regime and decreases in the High EPU Regime. A possible explanation for the decrease in the EPUI in the High EPU Regime could be that under already heightened EPU, an unexpected increase in CPU may induce a shift in focus from broader economic policy issues to climate policy.
Consistent with the previous literature (Fried et al., 2022, Khalil and Strobel, 2023, and Huang and Punzi, 2024), a positive CPU shock decreases production in both regimes. Huang and Punzi (2024) attribute the decrease in production to firms postponing investments, whereas Fried et al. (2022) and Khalil and Strobel (2023) consider mechanisms that shift investments toward cleaner capital. Although many of the GIRFs exhibit a short-term increase of IPI in the High EPU Regime, the negative long-term response of IPI seems, on average, roughly similar in both of the regimes.
In both regimes, a positive CPU shock increases consumer prices, in line with Huang and Punzi (2024) and Khalil and Strobel (2023). Huang and Punzi (2024) found that a positive CPU shock increases inflation by acting as a negative supply shock, whereas according to Khalil and Strobel (2023), a positive CPU shock may decrease aggregate demand due to risk-averse households raising savings, while firms raise their prices pre-emptively to lower the probability of being stuck with a negative markup. However, we find that the inflationary effects are stronger in the High EPU Regime than in the Low EPU Regime. One possible explanation is that the firms are more risk averse in the periods of high EPU, and thereby react more sensitively. Producer prices, in turn, mostly decrease in the Low EPU Regime but increase in the High EPU Regime.
The response of the interest rate variable seems consistent with the inflationary effects of the CPU shock: it increases in both regimes, but clearly more so in the High EPU Regime. Finally, consistent with the GIRFs of EPUI, a positive CPU shock temporarily increases the transition weights of the High EPU Regime in the periods of low EPU, while the effect is opposite in the periods of high EPU. Overall, our results suggest that a positive CPU shock decreases production and increases inflation in both regimes, but its inflationary effects are stronger in the periods of high EPU than in times of low EPU.
6 Conclusion
Linear structural vector autoregressive models can be identified statistically by non-Gaussianity without imposing restrictions on the model if the shocks are mutually independent and at most one of them is Gaussian (Lanne et al., 2017). We have shown that this result extends to structural threshold and smooth transition vector autoregressive models incorporating a time-varying impact matrix defined as a weighted sum of the impact matrices of the regimes. The weights of the impact matrices are the transition weights of the regimes, which we assume to be either exogenous, logistic, or of the threshold form, but our results can be generalized to other suitable weight functions as well. Thus, the types of transition weights typically used in macroeconomic applications are accommodated.
In addition to establishing the identification of the shocks, we have discussed the problem of labelling them, and employed blended identification (cf. Carriero et al., 2024) to address weak identification with respect to the ordering and signs of the columns of the regime-specific impact matrices. We have proposed estimating the model parameters by the method of penalized maximum likelihood, and introduced a three-step estimation procedure by adapting the approach of Lanne et al. (2017) to nonlinear SVAR models. Building on the results of Saikkonen (2008), we have also provided a sufficient condition for ergodic stationarity of our structural STVAR model. The introduced methods have been implemented in the accompanying R package sstvars (Virolainen, 2025a), which is available via the CRAN repository.
In an empirical application to U.S. data from 1987:4 to 2024:12, we have studied the macroeconomic effects of the climate policy uncertainty shock using a two-regime logistic STVAR model. As a measure of climate policy uncertainty, we employed the CPU index (Gavriilidis, 2021) constructed based on the amount of newspaper coverage on climate policy uncertainty topics. To distinguish between periods low and high economic policy uncertainty, we used the first lag of the EPU index (Baker et al., 2016) as the switching variable. Our results are in line with the previous literature suggesting that a positive CPU shock decreases production (Fried et al., 2022, Khalil and Strobel, 2023, and Huang and Punzi, 2024) and increases inflation (Khalil and Strobel, 2023, and Huang and Punzi, 2024). However, we found that while a positive CPU shock decreases production and increases inflation in times of both low and high EPU, its inflationary effects are stronger during the periods of high EPU.
Appendix A Stationarity and ergodicity
Establishing stationarity and ergodicity of the model is a common practice in time series econometrics, as it facilitates obtaining desirable asymptotic properties such as consistency of the estimator. Saikkonen (2008) derives a sufficient condition for ergodic stationarity of the smooth transition vector error correction (STVEC) model. Kheifets and Saikkonen (2020) make use of the results of Saikkonen (2008) to show that the stationarity condition readily applies for a STVAR model that can be obtained as a special case of the STVEC model of Saikkonen (2008). However, in both Saikkonen (2008) and Kheifets and Saikkonen (2020), the conditional covariance matrix of the process is defined as a weighted sum of positive definite error term covariance matrices of the regimes. As discussed in Section 3.1, the conditional covariance matrix (3.4) of our structural STVAR model has a different form, and therefore, our model is not nested to the model of Saikkonen (2008). The applicability of the results of Saikkonen (2008) also requires that the weights are not exogenous but functions of the preceding $p$ observations (the time $t$ weights are allowed to depend on the time $t$ observation of an IID process independent of $e_{t}$, but we do not discuss this option further, as it violates our assumption of $\mathcal{F}_{t-1}$-measurable or nonrandom weights). Nonetheless, we will show in this section that the stationarity condition applies to our model as well.
The sufficient stationarity condition is expressed in terms of the joint spectral radius (JSR) of certain matrices. The JSR of a finite set of square matrices $\mathcal{A}$ is defined by
$\rho(\mathcal{A})=\underset{j\rightarrow\infty}{\limsup}\left(\underset{A\in\mathcal{A}^{j}}{\sup}\rho(A)\right)^{1/j},$ (A.1)
where $\mathcal{A}^{j}=\{A_{1}A_{2}...A_{j}:A_{i}\in\mathcal{A}\}$ and $\rho(A)$ is the spectral radius of the square matrix $A$.
Consider the companion form AR matrices of the regimes defined as
$\boldsymbol{A}_{m}=\underset{(dp\times dp)}{\begin{bmatrix}A_{m,1}&A_{m,2}&\cdots&A_{m,p-1}&A_{m,p}\\ I_{d}&0&\cdots&0&0\\ 0&I_{d}&&0&0\\ \vdots&&\ddots&\vdots&\vdots\\ 0&0&\ldots&I_{d}&0\end{bmatrix}},\ m=1,...,M.$ (A.2)
The following assumption collects the sufficient conditions for establishing ergodicity, stationarity, and mixing properties of our structural STVAR model.
Assumption A.1.
Suppose the following conditions hold:
  1. (I)
    $\rho(\{\boldsymbol{A}_{1},...,\boldsymbol{A}_{M}\})<1$,
  2. (II)
    The distribution of the $IID(0,I_{d},\nu)$ random vector $e_{t}$ has a (Lebesgue) density that is bounded away from zero on compact subsets of $\mathbb{R}^{d}$.
  3. (III)
    The transition weight $\alpha_{m,t}$ is an $\mathcal{F}_{t-1}$-measurable function of $\{y_{t-j},j=1,...,p\}$ for all $t$ and $m=1,...,M$.
Condition (I) states that the JSR of the companion form AR matrices of the regimes is strictly less than one, and it is analogous to Condition (10) of Saikkonen (2008). Note that this condition is sufficient but not necessarily necessary for ergodic stationarity of the process (to the best of our knowledge, there does not currently exist results presenting a sufficient necessary condition for STVAR models). Condition (II) is analogous Assumption 1 of Saikkonen (2008), and it is innocuous in practice but rules out bounded error distributions, for example. Finally, Condition (III) ensures that the assumptions for the transition function imposed by Saikkonen (2008) are satisfied, ruling out nonrandom exogenous weights (which are accommodated by our identification results).
The following theorem, which is analogous to Theorem 1 in Saikkonen (2008) and Theorem 1 in Kheifets and Saikkonen (2020), states the results.
Theorem A.1.
Consider the STVAR process $y_{t}$ defined in Equations (3.2) and (3.3), and suppose Assumptions 1 and A.1 are satisfied. Then, almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, the process $\boldsymbol{y}_{t}=(y_{t},...,y_{t-p+1})$ is a $(1+||x||^{2})$-geometrically ergodic Markov chain. Thus, there exists a choice of initial values $y_{-p+1},...,y_{0}$ such that the process $y_{t}$ is strictly stationary, second-order stationary, and $\beta$-mixing with geometrically decaying mixing numbers.
Theorem A.1 is proven in Appendix A.1. A notable difference of Theorem A.1 to Theorem 1 of Saikkonen (2008) (and Theorem 1 of Kheifets and Saikkonen, 2020) is that our results holds only almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, while the result of Saikkonen (2008) holds everywhere. To be more specific, the result of Theorem A.1 holds with such values of $B_{1},...,B_{M}$ that the resulting impact matrix $B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ is invertible for all $t$, and Lemma 1 shows that this holds everywhere expect in regions of $\mathbb{R}^{d\times dM}$ that have a Lebesgue measure zero.
Assumption A.1 is sufficient for ergodic stationarity of our structural STVAR model, but Assumption A.1(I) is computationally demanding to verify in practice (e.g., Chang and Blondel, 2013). This makes it poorly suitable for restricting the parameter space in the numerical estimation of parameters discussed in Section 4. During estimation, this condition would have to be verified repeatedly a very large of times, making the estimation extremely tedious. Hence, it is useful make use of the following condition in the estimation, which is necessary for Assumption A.1(I).
Condition A.1.
$\max\{\rho(\boldsymbol{A}_{1}),...,\rho(\boldsymbol{A}_{M})\}<1$,
where $\rho(\boldsymbol{A}_{m})$ is the spectral radius of $\boldsymbol{A}_{m}$, $m=1,...,M$.
Condition A.1 states that the usual stability condition is satisfied by each of the regimes. It is necessary for Assumption A.1(I), as $\max(\rho(\boldsymbol{A}_{1}),...,\rho(\boldsymbol{A}_{M}))\leq\rho(\{\boldsymbol{A}_{1},...,\boldsymbol{A}_{M}\})$ (see Kheifets and Saikkonen, 2020, and the references therein). After estimation, Assumption A.1(I) can be checked for the solutions of interest.
Several methods have been proposed for bounding the JSR, many of which are discussed in Chang and Blondel (2013). The accompanying R package sstvars (Virolainen, 2025a) implements the branch-and-bound method of Gripenberg (1996), but the implementation is computationally very demanding if the matrices are large and a tight bound is required. The JSR toolbox in MATLAB (Jungers, 2023), in turn, automatically combines various methods to substantially enhance computational efficiency.
Appendix B Proofs
B.1 Proof of Lemma 1
Before proving Lemma 1, we establish the following lemma that is made use of in the proof:
Lemma B.1.
Suppose $P(x_{1},...,x_{k})$ is a finite-degree polynomial in the scalar variables $x_{i}\in\mathbb{R}$, $i=1,...,k$. Denote the set of vectors $(x_{1},...,x_{k})$ that satisfy the equation
$P(x_{1},...,x_{k})=0$ (B.1)
as $\mathcal{X}\equiv\{(x_{1},...,x_{k})\in\mathbb{R}^{k}|P(x_{1},...,x_{k})=0\}$. Then, the set $\mathcal{X}$ has a Lebesgue measure zero in $\mathbb{R}^{k}$, unless Equation (B.1) is identically true.
Proof.
If $k=1$, the proof is trivial. Suppose $k>1$. It then follows from Implicit function theorem that in the points $x^{0}=(x_{1,0},...,x_{k,0})\in\mathbb{R}^{k}$ where the gradient of $P(x_{1},...,x_{k})$ is nonzero, the set $\mathcal{X}$ is an $k-1$-dimensional manifold locally in a neighbourhood of $x^{0}$. Because this argument applies at all points $x^{0}$ where the gradient of $P(x_{1},...,x_{k})$ is nonzero, it follows that the set $\mathcal{X}$ is an $k-1$-dimensional manifold in regions where this condition holds. Therefore, Lebesque measure of the set $\mathcal{X}$ is zero in $\mathbb{R}^{k}$ in the region where the gradient of $P(x_{1},...,x_{k})$ is nonzero. On the other hand, by Sard’s theorem (Sard, 1994), the set of points where the gradient of $P(x_{1},...,x_{k})$ is zero has a Lebesgue measure zero in $\mathbb{R}^{k}$. Since the union of two sets with a Lebesgue measure zero has a Lebesgue measure zero, these results can be combined to conclude that the $\mathcal{X}$ has a Lebesgue measure zero in $\mathbb{R}^{k}$. ∎
To prove Lemma 1, we start by showing that for all $\alpha_{1,t},...,\alpha_{M,t}$, the matrix $B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ is invertible almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$. First, note that by the constraints $\sum_{m=1}\alpha_{m,t}=1$ and $\alpha_{m,t}\geq 0$, $m=1,...,M$, we have $\alpha_{m,t}>0$ for some $m$ for all $t$. Therefore, for all $\alpha_{1,t},...,\alpha_{M,t}$, there exists invertible matrices $B_{1},...,B_{M}$ such that $B_{y,t}$ is invertible. The set of points $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$ (or equally $(\text{vec}(B_{1}),...,\text{vec}(B_{M}))\in\mathbb{R}^{Md^{2}}$) for which $B_{y,t}$ is singular is $S_{t}\equiv\{(\text{vec}(B_{1}),...,\text{vec}(B_{M}))\in\mathbb{R}^{Md^{2}}|\det(\sum_{m=1}^{M}\alpha_{m,t}B_{m})=0\}$. Since the determinant of a matrix is polynomial in the entries of that matrix, and the entries of $\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ are linear combinations of the entries of $B_{1},...,B_{M}$, it follows that the determinant of $\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ is a polynomial in the entries of $B_{1},...,B_{M}$. Denote this polynomial as $P_{t}(B_{1},...,B_{M})\equiv\det(\sum_{m=1}^{M}\alpha_{m,t}B_{m})$. Note that whenever $\alpha_{m,t}=0$, the impact matrices $B_{m}$ of the corresponding regimes drop out from $P_{t}(B_{1},...,B_{M})$, but in this case $P_{t}(B_{1},...,B_{M})$ can still be interpreted as a polynomial in the entries of $B_{1},...,B_{M}$ with the multipliers of the dropped entries being zero. It then follows from Lemma B.1 that $S_{t}$ has a Lebesgue measure zero in $\mathbb{R}^{Md^{2}}$.
Finally, note that since the above result holds for all $\alpha_{1,t},...,\alpha_{M,t}$, the set $S_{t}$ has a Lebesgue measure zero in $\mathbb{R}^{Md^{2}}$ for each $t$. Hence, the union $\cup_{t=1}^{\infty}S_{t}$ is a union of a countable number of measure zero sets and thus itself has a Lebesgue measure zero. It follows that the matrix $B_{y,t}$ is invertible for all $t$ almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, concluding the proof.∎
B.2 Proof of Lemma 2
The first part of our proof is largely similar to the proof of Proposition 1 in Lanne et al. (2017) building on results from independent component analysis (Comon, 1994). First, note that from Equation (3.2), we can rewrite the model as
$y_{t}-\phi_{y,t}-\boldsymbol{A}_{y,t}\boldsymbol{y}_{t-1}=B_{y,t}e_{t},$ (B.2)
where $\phi_{y,t}\equiv\sum_{m=1}^{M}\alpha_{m,t}\phi_{m}$ $(d\times 1)$, $\boldsymbol{A}_{y,t}\equiv[A_{y,t,1}:...:A_{y,t,p}]$ $(d\times dp)$, and $A_{y,t,i}\equiv\sum_{m=1}^{M}\alpha_{m,t}A_{m,i}$ contain the time-varying AR coefficients, and $\boldsymbol{y}_{t-1}=(y_{t-1},...,y_{t-p})$. Then, suppose Equation (B.2) holds also for some other coefficients, transition weights, impact matrix, and shocks $\tilde{\phi}_{m},\tilde{A}_{m,1},...,\tilde{A}_{m,p},\tilde{\alpha}_{m,t}$, $m=1,...,M$, $\tilde{B}_{y,t}$, and $\tilde{e}_{t}$ (and define $\tilde{\phi}_{y,t}$ and $\tilde{\boldsymbol{A}}_{y,t}$ analogously), i.e., $y_{t}-\tilde{\phi}_{y,t}-\tilde{\boldsymbol{A}}_{y,t}\boldsymbol{y}_{t-1}=\tilde{B}_{y,t}\tilde{e}_{t}$. By subtracting this identify from (B.2), we obtain
$\tilde{\phi}_{y,t}-\phi_{y,t}+(\tilde{\boldsymbol{A}}_{y,t}-\boldsymbol{A}_{y,t})\boldsymbol{y}_{t-1}=B_{y,t}e_{t}-\tilde{B}_{y,t}\tilde{e}_{t}.$ (B.3)
Taking conditional expectation conditional on $\mathcal{F}_{t-1}$ gives
$\tilde{\phi}_{y,t}-\phi_{y,t}+(\tilde{\boldsymbol{A}}_{y,t}-\boldsymbol{A}_{y,t})\boldsymbol{y}_{t-1}=0,$ (B.4)
as $E[B_{y,t}e_{t}|\mathcal{F}_{t-1}]=E[\tilde{B}_{y,t}\tilde{e}_{t}|\mathcal{F}_{t-1}]=0$ (by $e_{t}\perp\!\!\!\!\perp\mathcal{F}_{t-1}$ and $E[e_{t}]=0$). By substituting this back to (B.3), we get
$B_{y,t}e_{t}=\tilde{B}_{y,t}\tilde{e}_{t}.$ (B.5)
Since $B_{y,t}$ is invertible almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$ by Lemma 1, we can solve $e_{t}$ as $e_{t}=C\tilde{e}_{t}$, where $C=B_{y,t}^{-1}\tilde{B}_{y,t}$ is $\mathcal{F}_{t-1}$-measurable. By Assumption 1, the random variables $\tilde{e}_{1t},...,\tilde{e}_{dt}$ are mutually independent and at most one of them is Gaussian. Also the random variables $e_{1t},...,e_{dt}$ are mutually independent and at most one of them is Gaussian. Therefore, conditionally on $\mathcal{F}_{t-1}$, by Darmois-Skitovich theorem (Lemma A.1 in Lanne et al., 2017), at most one column of $C$ may contain more than one nonzero element. Suppose, say, the $k$th column of $C$ has at least two nonzero elements, $c_{ik}$ and $c_{jk}$, $i\neq j$. Then, $e_{it}=c_{ik}\tilde{e}_{kt}+\sum_{l=1,l\neq k}^{d}c_{il}\tilde{e}_{lt}$ and $e_{jt}=c_{jk}\tilde{e}_{kt}+\sum_{l=1,l\neq k}^{d}c_{jl}\tilde{e}_{lt}$ with $\tilde{e}_{kt}$ Gaussian (by Darmois-Skitovich theorem). This implies that $E[e_{it}e_{jt}]=c_{ik}c_{jk}\neq 0$ (since $E[\tilde{e}_{kt}^{2}]=1$), so $e_{it}$ and $e_{jt}$ are not independent, which is a contradiction. Hence, each column of $C$ has at most one nonzero element.
Now, from the invertibility of $C$ it follows that each column of $C$ as exactly one nonzero element, and for the same reason, also that each row of $C$ has exactly one nonzero element. Therefore, there exists a permutation matrix $P$ and a diagonal matrix $D=\text{diag}(d_{1},...,d_{d})$ with nonzero diagonal elements such that $C=DP$. Together with $C=B_{y,t}^{-1}\tilde{B}_{y,t}$ this implies that $\tilde{B}_{y,t}=B_{y,t}DP$ and $\tilde{e}_{t}=P^{\prime}D^{-1}e_{t}$, concluding that $B_{y,t}$ is unique up to ordering of its columns and multiplying each column by a constant.
To complete the first part of the proof, it needs to be shown that the diagonal elements of $D$ are either $1$ or $-1$. By the normalization of the unit variances of $e_{t}$ and $\tilde{e}_{t}$, $\text{Cov}(\tilde{e}_{t}|\mathcal{F}_{t-1})=\text{Cov}(Ce_{t}|\mathcal{F}_{t-1})=I_{d}$, which is equivalent to $C\text{Cov}(e_{t}|\mathcal{F}_{t-1})C^{\prime}=I_{d}$ and thereby $CC^{\prime}=I_{d}$. By substituting $C=DP$ to this expression, we obtain $DPP^{\prime}D^{\prime}=I_{d}$ and by the orthogonality of permutation matrix $DD^{\prime}=I_{d}$. That is, $d_{ii}^{2}=1$ for all $i=1,...,d$, so $d_{ii}=\pm 1$. Since diagonal elements of $D$ are, hence, either $1$ or $-1$, it follows that $B_{y,t}$ is unique up to ordering of its columns and changing all signs in a column almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.
Threshold weights. To show the ”moreover”-part, consider first the threshold transition weights defined in (2.3). Denote the partition of the time periods to the regimes as $\mathcal{T}_{m}=\{t:r_{m-1}<z_{t}\leq r_{m}\}$, $m=1,...,M$, where $z_{t}$ is the switching variable and $r_{0}=-\infty,r_{M}=\infty$, and $r_{1},...,r_{M-1}$ are threshold parameters, and suppose Equation (B.2) holds also for some other partition $\tilde{\mathcal{T}}_{m}=\{t:\tilde{r}_{m-1}<z_{t}\leq\tilde{r}_{m}\}$, $m=1,...,M$. By Condition (3) of Lemma 2, the sets $\mathcal{T}_{m}$ and $\tilde{\mathcal{T}}_{m}$ are nonempty for all $m=1,...,M$. That is, we assume Equation (B.4) holds for the two sets of parameters $\phi_{m},A_{m,1},...,A_{m,p}$, $m=1,...,M$, $r_{1},...,r_{M-1}$ and $\tilde{\phi}_{m},\tilde{A}_{m,1},...,\tilde{A}_{m,p}$, $m=1,...,M$, $\tilde{r}_{1},...,\tilde{r}_{M-1}$.
At each $t$, the model reduces to a linear VAR $y_{t}=\phi_{m}+\sum_{i=1}^{p}A_{m,i}y_{t-1}+B_{m}e_{t}$ corresponding to the regime $m$ for which $r_{m-1}<z_{t}\leq r_{m}$. Consequently, Equation (B.4) can be written as
$\tilde{\phi}_{\tilde{m}_{t}}-\phi_{m_{t}}+\sum_{i=1}^{p}(\tilde{A}_{\tilde{m}_{t},i}-A_{m_{t},i})y_{t-i}=0,$ (B.6)
where $m_{t}\in\{1,...,M\}$ is the regime $m$ that satisfies $r_{m-1}<z_{t}\leq r_{m}$ and $\tilde{m}_{t}\in\{1,...,M\}$ is the regime $m$ that satisfies $\tilde{r}_{m-1}<z_{t}\leq\tilde{r}_{m}$ at the time $t$.
Equation (B.6) defines $d$ scalar equations such that the left side of each scalar equation is a first degree polynomial in $\boldsymbol{y}_{t-1}\in\mathbb{R}^{dp}$, where, for the $j$th equation, the coefficients are given by the $j$th element of $\tilde{\phi}_{\tilde{m}_{t}}-\phi_{m_{t}}$ and the elements of the $j$th rows of $\tilde{A}_{\tilde{m}_{t},i}-A_{m_{t},i}$, $i=1,...,p$. Hence, by applying Lemma B.1 to each scalar equation in (B.6), it can be concluded that the set of solutions to Equation (B.6) has a Lebesgue measure zero in $\mathbb{R}^{dp}$ unless it is identically true. Since Equation (B.6) is identically true if and only if $\tilde{\phi}_{\tilde{m}_{t}}=\phi_{m_{t}}$ and $\tilde{A}_{\tilde{m}_{t},i}=A_{m_{t},i}$, $i=1,...,p$, and because Equation (B.6) must hold for all $t$ (thereby for all regimes), it follows that $\tilde{\phi}_{\tilde{m}}=\phi_{m}$ and $\tilde{A}_{\tilde{m},i}=A_{m,i}$, $i=1,...,p$, for all $m=1,...,M$ almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ (as a countable union of measure zero sets has a measure zero). It then follows from Condition (2) of Lemma 2 that once $\tilde{\phi}_{\tilde{m}}=\phi_{m}$ and $\tilde{A}_{\tilde{m},i}=A_{m,i}$, $i=1,...,p$, for all $m=1,...,M$, the only way for these parameters to match the data almost everywhere is that the same regime index must apply for each $t$. Consequently, $\tilde{m}_{t}=m_{t}$ for all $t$, so the threshold parameters $r_{1},...,r_{M-1}$ and $\tilde{r}_{1},...,\tilde{r}_{M-1}$ imply identical partitions of the data to the regimes, $\mathcal{\tilde{T}}_{m}=\mathcal{T}_{m}$ for all $m=1,...,M$ almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$.
Because a regime switches when (and only when) the switching variable passes one of the threshold parameter values, for a given (countable number of) observations of $z_{t}$, the threshold parameters can only be identified with the resolution determined by the spacing of the order statistics of $z_{t}$. For instance, if the $n$th ($n\in\{1,...,T-1\}$) order statistic of $z_{t}$ is $1.00$, its $n+1$th order statistic is $1.02$, and $M=2$, then all values of the threshold parameter $r_{1}$ between $1.00$ and $1.02$ correspond to the same model, as they all imply the same regime-switches. Hence, the identification of the threshold parameters should be understood in the sense that they are identified to be between some order statistics of the switching variable series $z_{t}$. Now, if the threshold parameter values $\tilde{r}_{1},...,\tilde{r}_{M-1}$ are between different order statistics of $z_{t}$ than $r_{1},...,r_{M-1}$, the partitions $\mathcal{\tilde{T}}_{m}$ and $\mathcal{T}_{m}$ are different for some $m\in\{1,...,M\}$, in which case, by the above discussion, Equation (B.6) can only hold in sets of Lebesgue measure zero in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$. This concludes that the parameters $\phi_{m},A_{m,1},...,A_{m,p}$, $m=1,...,M$ and $r_{1},...,r_{M-1}$ are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$.
Logistic weights. Next, we show the ”moreover” part for the logistic transition weights defined in (2.4) with $M=2$. Suppose Equation (B.4) holds for the two sets of parameters $\phi_{m},A_{m,1},...,A_{m,p}$, $m=1,...,M$, $c,\gamma$ and $\tilde{\phi}_{m},\tilde{A}_{m,1},...,\tilde{A}_{m,p}$, $m=1,...,M$, $\tilde{c},\tilde{\gamma}$. The proof proceeds in the following steps: (1) we show that $\tilde{\alpha}_{m,t}=\alpha_{m,t}$ for each $t$ almost everywhere in $z_{t}\in\mathbb{R}$, (2) we use this identity to show that the parameters $c,\gamma$ are uniquely identified (almost everywhere), and then (3) we use these results to show that the intercepts and autoregression matrices are uniquely identified.
By applying Lemma B.1 to each scalar equation in Equation (B.4), it follows that Equation (B.4) holds almost everywhere in $\boldsymbol{y}_{t-1}\in\mathbb{R}^{dp}$ if only if Equation (B.4) is identically true, that is, $\tilde{\phi}_{y,t}=\phi_{y,t}$ and $\tilde{\boldsymbol{A}}_{y,t}=\boldsymbol{A}_{y,t}$ (irrespective of the weight function dynamics). By making use of the identity $\alpha_{1,t}=1-\alpha_{2,t}$, these equations can be written as
$\tilde{\phi}_{1}-\phi_{1}+\alpha_{2,t}(\phi_{1}-\phi_{2})-\tilde{\alpha}_{2,t}(\tilde{\phi}_{1}-\tilde{\phi}_{2})=0$ (B.7)
and
$\tilde{A}_{1,i}-A_{1,i}+\alpha_{2,t}(A_{1,i}-A_{2,i})-\tilde{\alpha}_{2,t}(\tilde{A}_{1,i}-\tilde{A}_{2,i})=0,\ \ i=1,...,p,$ (B.8)
respectively.
Equations (B.7) and (B.8) define a system of $d(1+pd)$, $d\geq 2,p\geq 1$, linear scalar equations in two variables $\alpha_{2,t}$ and $\tilde{\alpha}_{2,t}$ (interpreting the AR parameters as coefficients here). Suppose first that all of them are identically true, i.e., that all the coefficients are zero. In (B.7), this requires that $\tilde{\phi}_{1}=\phi_{1}$, $\phi_{1}=\phi_{2}$, and $\tilde{\phi}_{1}=\tilde{\phi}_{2}$. Similarly in (B.8), zero coefficients require that $\tilde{A}_{1,i}=A_{1,i}$, $A_{1,i}=A_{2,i}$ and $\tilde{A}_{1,i}=\tilde{A}_{2,i}$ for all $i=1,...,p$. But as $\phi_{1}=\phi_{2}$ and $A_{1,i}=A_{2,i}$ for all $i=1,...,p$ violates Condition (2) of Lemma 2, it must be that not all scalar equations in (B.7) and (B.8) are identically true.
For each scalar equation that is not identically true, the coefficients are not all zero, and each of these scalar equations can be expressed in the form $a_{2,j}\alpha_{2,t}=a_{0,j}+a_{3,j}\tilde{\alpha}_{2,t}$ for some $a_{0,j},a_{2,j},a_{3,j}\in\mathbb{R}$ for the $j$th not identically true equation. If $a_{2,j}=0$, then $\tilde{\alpha}_{2,t}$ is the constant $-a_{0,j}/a_{3,j}$ for all $t$, which violates Condition (3) of Lemma 2. Hence, $a_{2,j}\neq 0$ (in the not identically true equations), implying that $\alpha_{2,t}=a_{0,j}+a_{1,j}\tilde{\alpha}_{2,t}$, where $a_{1,j}\equiv a_{3,j}a_{2,j}^{-1}$. Thus, the solution to Equations (B.7) and (B.8) must satisfy this relationship (for all $j$).
Making the dependence of the transition weights of $z_{t}$ explicit in the notation, suppose $\alpha_{2,t}(z_{t})=a_{0,j}+a_{1,j}\tilde{\alpha}_{2,t}(z_{t})$ holds in a set of positive Lebesque measure in $z_{t}\in\mathbb{R}$. Since $\alpha_{2,t}(z_{t})$ and $\tilde{\alpha}_{2,t}(z_{t})$ are logistic, $\alpha_{2,t}(z_{t})$ and $a_{0,j}+a_{1,j}\tilde{\alpha}_{2,t}(z_{t})$ are real analytic functions. Thus, by the Identity theorem, we have that $\alpha_{2,t}(z_{t})=a_{0,j}+a_{1,j}\tilde{\alpha}_{2,t}(z_{t})$ for all $z_{t}$. Since this must hold for all $z_{t}$, it follows from $\underset{z_{t}\rightarrow-\infty}{\lim}\alpha_{2,t}(z_{t})=0$ and $\underset{z_{t}\rightarrow-\infty}{\lim}(a_{0,j}+a_{1,j}\tilde{\alpha}_{2,t}(z_{t}))=a_{0,j}$ that $a_{0,j}=0$. Similarly, assuming $a_{1,j}\neq 0$, it follows from $a_{0,j}=0$, $\underset{z_{t}\rightarrow+\infty}{\lim}\alpha_{2,t}(z_{t})=1$, and $\underset{z_{t}\rightarrow+\infty}{\lim}(a_{0,j}+a_{1,j}\tilde{\alpha}_{2,t}(z_{t}))=a_{0,j}+a_{1,j}$ that $a_{1,j}=1$. If $a_{1,j}=0$, then clearly $\alpha_{2,t}(z_{t})=a_{0,j}+a_{1,j}\tilde{\alpha}_{2,t}(z_{t})$ holds only in the limit $z_{t}\rightarrow-\infty$, and thereby not in a set of positive Lebesgue measure. Thus, it must be that $a_{0,j}=0$ and $a_{1,j}=1$ in sets of positive Lebesgue measure, concluding that $\alpha_{2,t}(z_{t})=\tilde{\alpha}_{2,t}(z_{t})$ almost everywhere in $z_{t}\in\mathbb{R}$.
Next, we make use of the identity $\alpha_{2,t}(z_{t})=\tilde{\alpha}_{2,t}(z_{t})$ to show that $\tilde{c}=c$ and $\tilde{\gamma}=\gamma$. By substituting the logistic weight function and simplifying, the identity $\alpha_{2,t}(z_{t})=\tilde{\alpha}_{2,t}(z_{t})$ can be expressed in the form
$(\tilde{\gamma}-\gamma)z_{t}=\tilde{\gamma}\tilde{c}-\gamma c.$ (B.9)
If $\tilde{\gamma}=\gamma$, it follows that $\tilde{c}=c$. If $\tilde{\gamma}\neq\gamma$, Equation (B.9) has only one solution, namely, $z_{t}=(\tilde{\gamma}-\gamma)^{-1}(\tilde{\gamma}\tilde{c}-\gamma c)$. Thus, $\tilde{\gamma}=\gamma$ and $\tilde{c}=c$ almost everywhere in $\boldsymbol{y}_{t-1}\in\mathbb{R}^{dp}$ and $z_{t}\in\mathbb{R}$.
We have shown that for each $t$, $\tilde{\gamma}=\gamma$ and $\tilde{c}=c$ holds almost everywhere in $\boldsymbol{y}_{t-1}\in\mathbb{R}^{dp}$ and $z_{t}\in\mathbb{R}$. If the $z_{t}$ is a function of $\boldsymbol{y}_{t-1}$ (rather than exogenous), we have that $\tilde{\gamma}=\gamma$ and $\tilde{c}=c$ almost everywhere in $\boldsymbol{y}_{t-1}\in\mathbb{R}^{dp}$ (as a countable union of measure zero sets has a measure zero). Since this holds for each $t$, combining it for $t=1,....,T$ yields that $\tilde{\gamma}=\gamma$ and $\tilde{c}=c$ almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ and $(z_{1},...,z_{T})\in\mathbb{R}^{T}$. If $z_{t}$ is a function of $\boldsymbol{y}_{t-1}$, this result holds almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$.
Before showing that the parameters $\phi_{m},A_{m,1},...,A_{m,p}$, $m=1,2$, are identified, we present the following lemma that is used in the proof:
Lemma B.2.
Consider the $T$ linear equations in $M$ variables
$\alpha_{1,t}x_{1}+...+\alpha_{M,t}x_{M}=c_{t},\quad t=1,...,T,$ (B.10)
where $x_{1},...,x_{M}$ are scalar variables, and $\alpha_{1,t},...,\alpha_{M,t}$ and $c_{t}$ are scalar constants. Suppose there exists $M$ indices $t_{1},...,t_{M}\in\{1,....,T\}$ such that the corresponding coefficient vectors $(\alpha_{1,t_{i}},...,\alpha_{M,t_{i}})$, $i=1,...,M$, are linearly independent. Then, there exists at most one solution to the system of linear equations (B.10).
Proof.
By assumption, there exists $M$ indices $t_{1},...,t_{M}$ such that the coefficient vectors $(\alpha_{1,t_{i}},...,\alpha_{M,t_{i}})$, $i=1,...,M$, are linearly independent. Now, the rows of the $(M\times M)$ matrix
$A^{\alpha}\equiv\begin{bmatrix}\alpha_{1,t_{1}}&\cdots&\alpha_{M,t_{1}}\\ \vdots&\ddots&\vdots\\ \alpha_{1,t_{M}}&\cdots&\alpha_{M,t_{M}}\\ \end{bmatrix}$ (B.11)
are linearly independent, and thus $A^{\alpha}$ is invertible. The $M$ linear equations corresponding to these indices form the system $A^{\alpha}x=c$, where $x=(x_{1},...,x_{M})$ and $c=(c_{t_{1}},...,c_{t_{M}})$. Because $A^{\alpha}$ is invertible, this system has exactly one solution.
Now, any solution $x$ of the full system (B.10) for $t=1,....,T$ must satisfy the above subsystem, which has exactly one solution. Therefore, the full system (B.10) has at most one solution. ∎
We have established that $\tilde{\gamma}=\gamma$ and $\tilde{c}=c$ (and $\alpha_{2,t}(z_{t})=\tilde{\alpha}_{2,t}(z_{t})$ for all $t$) almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ and, if the switching variable $z_{t}$ is exogenous, $(z_{1},...,z_{T})\in\mathbb{R}^{T}$. Using these identities, Equations (B.7) and (B.8) simplify to
$(1-\alpha_{2,t})\tilde{\phi}_{1}+\alpha_{2,t}\tilde{\phi}_{2}=(1-\alpha_{2,t})\phi_{1}+\alpha_{2,t}\phi_{2}$ (B.12)
and
$(1-\alpha_{2,t})\tilde{A}_{1,i}+\alpha_{2,t}\tilde{A}_{2,i}=(1-\alpha_{2,t})A_{1,i}+\alpha_{2,t}A_{2,i},\ \ i=1,...,p,$ (B.13)
respectively. Treating the parameters $\phi_{m},A_{m,1},...,A_{m,p}$, $m=1,2$, and $c,\gamma$, as well as ($\mathcal{F}_{t-1}$-measurable or nonrandom) transition weights $\alpha_{2,t}$ as coefficients here, at each $t$, each scalar equation in (B.12)-(B.13) is a linear equation in two variables given by the corresponding entries of $\tilde{\phi}_{m},\tilde{A}_{m,1},...,\tilde{A}_{m,p}$, $m=1,2$.
Each scalar equation in (B.12)-(B.13) defined together for all $t=1,...,T$ defines a system of $T$ linear equations and two variables. By Condition (3) of Lemma 2 (with $M=2$), there exists two indices $t_{1},t_{2}\in\{1,....,T\}$ such that the corresponding coefficient vectors $(\alpha_{1,t_{i}},\alpha_{2,t_{i}})$, or equally $(1-\alpha_{2,t_{i}},\alpha_{2,t_{i}})$, $i=1,2$, are linearly independent. Hence, it follows from Lemma B.2 that there exists at most one solution to each such system of scalar equations, and thereby also to the system of $Td(1+pd)$ linear equations (B.12)-(B.13), $t=1,....,T$, in $2d(1+pd)$ variables. Since one solution is $\tilde{\phi}_{m}=\phi_{m},\tilde{A}_{m,i}=A_{m,i}$, $i=1,...,p$, $m=1,2$, this is the unique solution. This concludes the proof that the parameters $\phi_{m},A_{m,1},...A_{m,p}$, $m=1,...,M$ and $c,\gamma$ are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ and, if the switching variable $z_{t}$ is exogenous, $(z_{1},...,z_{T})\in\mathbb{R}^{T}$.
Exogenous nonrandom weights. Suppose the transition weights $\alpha_{1,t},...,\alpha_{M,t}$ are exogenous (nonrandom). If the transition weights are binary, i.e., $\alpha_{m,t}\in\{0,1\}$ for all $m=1,...,M$ and $t=1,...,T$, then arguments similar to those used for threshold weights can be used to conclude that the parameters $\phi_{m},A_{m,1},...,A_{M,p}$, $m=1,...,M$, are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$.
Next, suppose the transition weights are not binary. By applying Lemma B.1 to each scalar equation in Equation (B.4), it follows that Equation (B.4) holds almost everywhere in $\boldsymbol{y}_{t-1}\in\mathbb{R}^{dp}$ if only if Equation (B.4) is identically true, that is, $\tilde{\phi}_{y,t}=\phi_{y,t}$ and $\tilde{\boldsymbol{A}}_{y,t}=\boldsymbol{A}_{y,t}$. Substituting the exogenous weights ($\alpha_{m,t}=\tilde{\alpha}_{m,t}$ is given for all $m=1,...,M$ and $t=1,...,T$) to these equations yields
$\sum_{m=1}^{M}\alpha_{m,t}(\tilde{\phi}_{m}-\phi_{m})$ $=0,\quad\text{and}$ (B.14)
$\sum_{m=1}^{M}\alpha_{m,t}(\tilde{A}_{m,i}-A_{m,i})$ $=0,\quad i=1,...,p,$ (B.15)
respectively.
Treating the parameters $\phi_{m},A_{m,1},...,A_{m,p}$, $m=1,2$, and $c,\gamma$, as well as transition weights $\alpha_{1,t},...,\alpha_{M,t}$ as constants, each scalar equations in (B.14)-(B.15) defined together for all $t=1,...,T$ defines a system of $T$ linear equations in $M$ variables. Therefore, arguments similar to those used for logistic weights but applied for any finite number of regimes regimes, making use of Lemma B.2, can be used to conclude that the parameters $\phi_{m},A_{m,1},...,A_{M,p}$, $m=1,...,M$, are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$.∎
B.3 Proof of Proposition 1
We start by showing that under the conditions of Proposition 1, the ordering and signs of the columns of $B_{y,t}$ are fixed for all $t$ in the sense that models before and after reordering or changing the signs of the columns of $B_{y,t}$ at any $t$ are not observationally equivalent. Specifically, we say that two models are observational equivalent if and only if their joint densities are equal. After that, we combine this result with Lemma 2 to show that $B_{1},...,B_{M}$ are uniquely identified.
First, we establish the following lemma and its corollary, the latter of which is used in the proof:
Lemma B.3.
Suppose $(\nu_{i},\lambda_{i})\neq(\nu_{j},\lambda_{j})$ (Condition (b) of Proposition 1). Then, $st(x_{1};\nu_{i},\lambda_{i})\neq st(x_{2};\nu_{j},\lambda_{j})$ almost everywhere in $(x_{1},x_{2})\in\mathbb{R}^{2}$, where the density function $st(\cdot;\nu_{i},\lambda_{i})$ is given in Equation (4.1). Moreover, if $\lambda_{i}\neq 0$, then the equation $st(x;\nu_{i},\lambda_{i})=st(-x;\nu_{i},\lambda_{i})$ has at most six solutions in $x\in\mathbb{R}$, and thus, $st(x;\nu_{i},\lambda_{i})\neq st(-x;\nu_{i},\lambda_{i})$ almost everywhere in $x\in\mathbb{R}$.
Proof.
Denote $f(x_{1},x_{2})\equiv st(x_{1};\nu_{i},\lambda_{i})-st(x_{2};\nu_{j},\lambda_{j})$, which is not identically zero by $(\nu_{i},\lambda_{i})\neq(\nu_{j},\lambda_{j})$. It is straightforward to show (by induction) that the function $st(x_{i};\nu_{i},\lambda_{i})$, and also thereby $f(x_{1},x_{2})$, is infinitely continuously differentiable. It then follows from Implicit function theorem that in the points $x^{0}=(x_{1,0},x_{2,0})\in\mathbb{R}^{2}$ where the gradient of $f(x_{1},x_{2})$ is nonzero, the set $\mathcal{X}\equiv\{(x_{1},x_{2})\in\mathbb{R}^{2}|f(x_{1},x_{2})=0\}$ is a one-dimensional manifold locally in a neighbourhood of $x^{0}$. Because this argument applies at all points $x^{0}$ where the gradient of $f(x_{1},x_{2})$ is nonzero, it follows that the set $\mathcal{X}$ is a one-dimensional manifold in regions where this condition holds. Therefore, Lebesque measure of the set $\mathcal{X}$ is zero in $\mathbb{R}^{2}$ in the region where the gradient of $f(x_{1},x_{2})$ is nonzero. On the other hand, by Sard’s theorem (Sard, 1994), the set of points where the gradient of $f(x_{1},x_{2})$ is zero has a Lebesgue measure zero in $\mathbb{R}^{2}$. Since the union of two sets with a Lebesgue measure zero has a Lebesgue measure zero, these results can be combined to conclude that the $\mathcal{X}$ has a Lebesgue measure zero in $\mathbb{R}^{2}$. That is, $st(x_{1};\nu_{i},\lambda_{i})=st(x_{2};\nu_{j},\lambda_{j})$ only on sets of zero Lebesgue measure, concluding that $st(x_{1};\nu_{i},\lambda_{i})\neq st(x_{2};\nu_{j},\lambda_{j})$ almost everywhere in $(x_{1},x_{2})\in\mathbb{R}^{2}$.
To prove the moreover part, suppose $\lambda_{i}\neq 0$ and $st(x;\nu_{i},\lambda_{i})=st(-x;\nu_{i},\lambda_{i})$, which simplifies to
$\left(\frac{b_{i}x+a_{i}}{1-\mathbbm{1}\{x<-a_{i}/b_{i}\}\lambda_{i}+\mathbbm{1}\{x\geq-a_{i}/b_{i}\}\lambda_{i}}\right)^{2}=\left(\frac{-b_{i}x+a_{i}}{1-\mathbbm{1}\{-x<-a_{i}/b_{i}\}\lambda_{i}+\mathbbm{1}\{-x\geq-a_{i}/b_{i}\}\lambda_{i}}\right)^{2},$ (B.16)
where $a_{i}\neq 0$ and $b_{i}\neq 1$. If $x<-a_{i}/b_{i}$ and $-x<-a_{i}/b_{i}$ or $x\geq-a_{i}/b_{i}$ and $-x\geq-a_{i}/b_{i}$, a direct calculation shows that the only solution to Equation (B.16) is $x=0$. If $x<-a_{i}/b_{i}$ and $-x\geq-a_{i}/b_{i}$ or $x\geq-a_{i}/b_{i}$ and $-x<-a_{i}/b_{i}$, Equation (B.16) is a second degree polynomial equation in $x$ and has thereby at most two real solutions. Combining all the four possible cases, in can be concluded that $st(x;\nu_{i},\lambda_{i})=st(-x;\nu_{i},\lambda_{i})$ has at most six solutions in $x\in\mathbb{R}$. ∎
Corollary B.1.
Suppose $D$ is a diagonal matrix with ±$\pm 1$ diagonal entries and $P$ is a permutation matrix, and denote by $I_{d,i}$ the $i$th column of the $d$-dimensional identity matrix. Moreover, suppose the entries $x$ are distinct as is and after changing any of their signs. If $(\nu_{i},\lambda_{i})\neq(\nu_{j},\lambda_{j})$ for all $i\neq j=1,...,d$ and $\lambda_{i}\neq 0$ for all $i=1,...,d$, and in addition, $D\neq I_{d}$ or $P\neq I_{d}$, then
$\prod_{i=1}^{d}st(I_{d,i}^{\prime}x;\nu_{i},\lambda_{i})\neq\prod_{i=1}^{d}st(I_{d,i}^{\prime}DPx;\nu_{i},\lambda_{i})$ (B.17)
almost everywhere in $x\in\mathbb{R}^{d}$.
Proof.
Suppose $\prod_{i=1}^{d}st(I_{d,i}^{\prime}x;\nu_{i},\lambda_{i})=\prod_{i=1}^{d}st(I_{d,i}^{\prime}DPx;\nu_{i},\lambda_{i})$. Rearranging and taking logarithm yields
$\sum_{i=1}^{d}(\log st(I_{d,i}^{\prime}x;\nu_{i},\lambda_{i})-\log st(I_{d,i}^{\prime}DPx;\nu_{i},\lambda_{i}))=0.$ (B.18)
Since $DPx$ reorders or swaps the signs of some of the elements of $x$ (but does not leave all the elements unchanged), by Lemma B.3 (and strict monotonicity of logarithm),
$f_{i}(x)\equiv\log st(I_{d,i}^{\prime}x;\nu_{i},\lambda_{i})-\log st(I_{d,i}^{\prime}DPx;\nu_{i},\lambda_{i})\neq 0$ (B.19)
for at least one $i\in\{1,...,d\}$ over sets of positive Lebesgue measure. Therefore, Equation (B.18) can hold almost everywhere only if the nonzero terms in the sum cancel each other out.
Denote $\mathcal{I}(x)\equiv\{i\in\{1,...,d\}|f_{i}(x)\neq 0\}$ and $g(x^{*})\equiv\sum_{i\in\mathcal{I}}f_{i}(x)$, where $x^{*}$ is the $(d^{*}\times 1)$-dimensional ($d^{*}\leq d$) subvector of $x$ containing the components of $x$ for which $I_{d,i}^{\prime}x\neq\pm I_{d,i}^{\prime}DPx$, $i=1,...,d$ (i.e., $x^{*}$ captures all variables contributing to any nonzero $f_{i}(x)$). It is straightforward to show (by induction) that the function $\log st(x_{i};\nu_{i},\lambda_{i})$, $x_{i}\in\mathbb{R}$, is infinitely continuously differentiable. It follows that also $g(x^{*})$ is infinitely continuously differentiable, and thus $d^{*}$ times continuously differentiable. Then, arguments based on Implicit function theorem and Sard’s theorem (Sard, 1994) can be used similarly to the proof of Lemma B.3 to conclude that the set of solutions to $g(x^{*})=0$ has a Lebesgue measure zero in $x^{*}\in\mathbb{R}^{d^{*}}$. Hence, as sets of zero Lebesgue measure in $\mathbb{R}^{d^{*}}\subset\mathbb{R}^{d}$ have zero Lebesgue measure in $\mathbb{R}^{d}$, Equation (B.18) can hold only in sets of zero Lebesgue measure in $x\in\mathbb{R}^{d}$, concluding the proof. ∎
Condition (a) of Proposition 1 fixes the ordering and signs of the columns of $B_{1}$, so it suffices to show that under Conditions (b) and (c), changing the ordering or signs of the columns of any of $B_{2},...,B_{M}$ would change the joint density of the model. Suppose the observed time series is $y_{t-p+1},...,y_{0},y_{1},...,y_{T}$ and denote $Y_{T}=(y_{1},...,y_{T})$. Conditional on the initial values $y_{t-p+1},...,y_{0}$, the joint density of the model is
$f_{Y}(Y_{T})=\prod_{t=1}^{T}\left(|\det(B_{y,t})|^{-1}\prod_{i=1}^{d}st(I_{d,i}^{\prime}B_{y,t}^{-1}(y_{t}-\mu_{y,t});\nu_{i},\lambda_{i})\right),$ (B.20)
where $I_{d,i}$ is the $i$th column of the $d$-dimensional identity matrix and $st(\cdot;\nu_{i},\lambda_{i})$ is the density function the $i$th shock given in Equation (4.1). At each $t$, $B_{y,t}=B_{m}$ for the regime $m$ with $r_{m-1}<z_{t}\leq r_{m}$, where the threshold parameters, as well as $B_{y,t}$ up to ordering and signs of its columns, are identified by Lemma 2 regardless of whether $B_{1},...,B_{M}$ are identified. Consequently the density (B.20) can be expressed as
$f_{Y}(Y_{T};B_{1},...,B_{M})$ $=\prod_{m=1}^{M}\prod_{t\in\mathcal{T}_{m}}f_{y}(y_{t}-\mu_{y,t};B_{m}),\ \ \text{and}$ (B.21)
$f_{y}(y_{t}-\mu_{y,t};B_{m})$ $=|\det(B_{m})|^{-1}\prod_{i=1}^{d}st(I_{d,i}^{\prime}B_{m}^{-1}(y_{t}-\mu_{y,t});\nu_{i},\lambda_{i}),$ (B.22)
where $\mathcal{T}_{m}$, $m=1,...,M$, is the set of time indices (identified by Lemma 2 almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ and nonempty by Condition (3) of Lemma 2) that satisfy $r_{m-1}<z_{t}\leq r_{m}$ for the regime $m$, $\mu_{y,t}$ is the conditional mean of the model (identified by Lemma 2 almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$), and the dependence of the density functions $f_{Y}(\cdot;B_{1},...,B_{M})$ and $f_{y}(\cdot;B_{m})$ on the impact matrices $B_{1},...,B_{M}$ is now made explicit in the notation (but their dependence on the other parameters is omitted).
Denote by $\boldsymbol{\tau}_{m}=(\tau_{m1},...,\tau_{md})$ the permutation $\tau_{m1},...,\tau_{md}$ of the columns of $B_{m}$, and denote by $\boldsymbol{\iota}_{m}=(\iota_{m1},...,\iota_{md})$ the sign changes $\iota_{m1},...,\iota_{md}$ of the columns of $B_{m}$. Moreover, denote by $B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})$ the impact matrix $B_{m}$ corresponding to the permutation $\boldsymbol{\tau}_{m}$ and sign changes $\boldsymbol{\iota}_{m}$ of its columns. Then, consider the permutations $\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{2},...,\boldsymbol{\tau}_{M}$ and $\boldsymbol{\tau}_{1},\tilde{\boldsymbol{\tau}}_{2},...,\tilde{\boldsymbol{\tau}}_{M}$ as well as the sign changes $\boldsymbol{\iota}_{1},\boldsymbol{\iota}_{2},...,\boldsymbol{\iota}_{M}$ and $\boldsymbol{\iota}_{1},\tilde{\boldsymbol{\iota}}_{2},...,\tilde{\boldsymbol{\iota}}_{M}$.
By the invertibility of $B_{m}$, almost everywhere in $y_{t}\in\mathbb{R}^{d}$, none of the entries in $B_{m}^{-1}(y_{t}-\mu_{y,t})$ are equal to each other as is nor after changing the signs of some of the entries. Since reordering the columns of $B_{m}$ reorders the rows of $B_{m}^{-1}$, it follows that if $\tilde{\boldsymbol{\tau}}_{m}\neq\boldsymbol{\tau}_{m}$, some of the entries in $B_{m}^{-1}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})(y_{t}-\mu_{y,t})$ and $B_{m}^{-1}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m})(y_{t}-\mu_{y,t})$ are in a different order. Likewise, changing the signs of the columns of $B_{m}$ changes the signs of the rows of $B_{m}^{-1}$, it follows that if $\tilde{\boldsymbol{\iota}}_{m}\neq\boldsymbol{\iota}_{m}$, some of the entries in $B_{m}^{-1}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})(y_{t}-\mu_{y,t})$ and $B_{m}^{-1}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m})(y_{t}-\mu_{y,t})$ have different signs (in addition to possibly being in a different order).
By Corollary B.1, different orderings and signs of the elements of $B_{m}^{-1}(y_{t}-\mu_{y,t})$ yield different values of $\prod_{i=1}^{d}st(I_{d,i}^{\prime}B_{m}^{-1}(y_{t}-\mu_{y,t});\nu_{i},\lambda_{i})$ almost everywhere in $y_{t}\in\mathbb{R}^{d}$ under Conditions (b) and (c) of Proposition 1. Moreover, as $|\det(B_{m})|$ is invariant to the ordering and signs of the columns of $B_{m}$, it follows from that, $f_{y}(y_{t}-\mu_{y,t};B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m}))\neq f_{y}(y_{t}-\mu_{y,t};B_{m}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m}))$ almost everywhere in $y_{t}\in\mathbb{R}^{d}$ whenever $\tilde{\boldsymbol{\tau}}_{m}\neq\boldsymbol{\tau}_{m}$ or $\tilde{\boldsymbol{\iota}}_{m}\neq\boldsymbol{\iota}_{m}$. Therefore, as $\mathcal{T}_{1},...,\mathcal{T}_{M}$ and $\mu_{y,t}$, $t=1,...,T$ are identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ by Lemma 2 and the shocks and thereby $y_{t}$ follow continuous distributions, we have
$\prod_{t\in\mathcal{T}_{m}}f_{y}(y_{t}-\mu_{y,t};B_{m}(\boldsymbol{\tau}_{m}))\neq\prod_{t\in\mathcal{T}_{m}}f_{y}(y_{t}-\mu_{y,t};B_{m}(\tilde{\boldsymbol{\tau}}_{m})),$ (B.23)
and thus also
$f_{Y}(Y_{T};B_{1}(\boldsymbol{\tau}_{1}),B_{2}(\boldsymbol{\tau}_{2}),...,B_{M}(\boldsymbol{\tau}_{M}))\neq f_{Y}(Y_{T};B_{1}(\boldsymbol{\tau}_{1}),B_{2}(\tilde{\boldsymbol{\tau}}_{2})...,B_{M}(\tilde{\boldsymbol{\tau}}_{M}))$ (B.24)
almost everywhere in $(y_{-p+1},...,y_{0},Y_{T})\in\mathbb{R}^{d(T+p)}$ whenever $\tilde{\boldsymbol{\tau}}_{m}\neq\boldsymbol{\tau}_{m}$ or $\tilde{\boldsymbol{\iota}}_{m}\neq\boldsymbol{\iota}_{m}$ for any $m=2,...,M$. Since this result holds almost everywhere in $(y_{-p+1},...,y_{0},Y_{T})\in\mathbb{R}^{d(T+p)}$, and $B_{y,t}$ is here invertible for all invertible $B_{1},...,B_{M}$, the result holds almost everywhere in $(y_{-p+1},...,y_{0},Y_{T})\in\mathbb{R}^{d(T+p)}$.
By Lemma 2, $B_{y,t}$ is identified up to ordering and signs of its columns at each $t$. Thus, each impact matrix $B_{1},...,B_{M}$ of the regimes is identified up to ordering signs of its columns. Under the conditions of Proposition 1, the ordering and signs of the columns of $B_{1}$ are directly fixed. Moreover, by the above discussion, different permutations or sign changes of the columns of any of $B_{2},...,B_{M}$ lead to different joint densities of the model almost everywhere in $(y_{-p+1},...,y_{0},Y_{T})\in\mathbb{R}^{d(T+p)}$, i.e., to models that are not observationally equivalent. Thus, only models corresponding to the same permutations and sign changes of the columns of $B_{2},...,B_{M}$ are observationally equivalent, implying that the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},Y_{T})\in\mathbb{R}^{d(T+p)}$.∎
B.4 Proof of Proposition 2
The proof consists of two parts. First, we show that if the ordering and signs of the columns of $B_{1}$ is fixed, changing the ordering or signs of the columns of any of $B_{2},...,B_{M}$ leads to an impact matrix $B_{y,t}$ that cannot, at all $t$, be obtained by reordering or changing the signs of the columns of the original impact matrix. In other words, we show that fixing the ordering signs of the columns of $B_{1}$ effectively fixes the ordering and signs of the columns of $B_{2},...,B_{M}$ and thereby of $B_{y,t}$. Then, we make use of this result together with the result of Lemma 2 (that $B_{y,t}$ is identified up ordering and signs of its columns at each $t$) to show that the matrices $B_{1},...,B_{M}$ are uniquely identified. For models with logistic transition weights (2.4), the result holds almost everywhere in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times dM}$, $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ and, if the switching variable $z_{t}$ is exogenous, $(z_{1},...,z_{T})\in\mathbb{R}^{T}$. For models with nonrandom exogenous transition weights, the result holds almost everywhere in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times dM}$.
By Lemma 2, for logistic models (2.4) (with $M=2$), the parameters $c,\gamma$, and threreby the transition weights $\alpha_{1,t},...,\alpha_{M,t}$$t=1,...,T$, are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},$ $y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ (and $(z_{1},...,z_{T})\in\mathbb{R}^{T}$ for exogenous $z_{t}$). With nonrandom exogenous transition weights, the identification of $\alpha_{1,t},...,\alpha_{M,t}$$t=1,...,T$ is given. In either case, the transition weights can be treated as uniquely identified in the rest of the proof, where, for readability, we do not repeatedly remind that the identification holds only almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ (and $(z_{1},...,z_{T})\in\mathbb{R}^{T}$ for exogenous $z_{t}$).
Denote by $\boldsymbol{\tau}_{m}=(\tau_{m1},...,\tau_{md})$ the permutation $\tau_{m1},...,\tau_{md}$ and $\boldsymbol{\iota}_{m}=(\iota_{m1},...,\iota_{md})$ the sign changes $\iota_{m1},...,\iota_{md}$ of the columns of $B_{m}$, respectively. Moreover, denote by $B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})$ the invertible impact matrix $B_{m}$ corresponding to the permutation $\boldsymbol{\tau}_{m}$ and sign changes $\boldsymbol{\iota}_{m}$ of its columns. Then, consider the permutations $\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{2},...,\boldsymbol{\tau}_{M}$ and $\boldsymbol{\tau}_{1},\tilde{\boldsymbol{\tau}}_{2},...,\tilde{\boldsymbol{\tau}}_{M}$ as well the sign changes $\boldsymbol{\iota}_{1},\boldsymbol{\iota}_{2},...,\boldsymbol{\iota}_{M}$ and $\boldsymbol{\iota}_{1},\tilde{\boldsymbol{\iota}}_{2},...,\tilde{\boldsymbol{\iota}}_{M}$. We start by showing that $B_{y,t}(\boldsymbol{\tau},\boldsymbol{\iota})\equiv\alpha_{1,t}B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})+\sum_{m=2}^{M}\alpha_{m,t}B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})$ cannot be obtained by reordering or changing the signs of the columns of $B_{y,t}(\tilde{\boldsymbol{\tau}},\tilde{\boldsymbol{\iota}})\equiv\alpha_{m,t}B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})+\sum_{m=2}^{M}\alpha_{m,t}B_{m}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m})$ almost everywhere in $[B_{1},...,B_{M}]\in\mathbb{R}^{d\times Md}$, whenever $\boldsymbol{\tau}_{m}\neq\tilde{\boldsymbol{\tau}}_{m}$ or $\boldsymbol{\iota}_{m}\neq\tilde{\boldsymbol{\iota}}_{m}$ for any $m=2,...,M$.
Suppose $\boldsymbol{\tau}_{m}\neq\tilde{\boldsymbol{\tau}}_{m}$ or $\boldsymbol{\iota}_{m}\neq\tilde{\boldsymbol{\iota}}_{m}$ for any $m=2,...,M$. At each time period $t$, the impact matrix $B_{y,t}(\tilde{\boldsymbol{\tau}},\tilde{\boldsymbol{\iota}})$ is obtained by reordering or changing the signs of the columns of $B_{y,t}(\boldsymbol{\tau},\boldsymbol{\iota})$ if and only if
$\alpha_{1,t}B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})+\sum_{m=2}^{M}\alpha_{m,t}B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})=(\alpha_{1,t}B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})+\sum_{m=2}^{M}\alpha_{m,t}B_{m}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m}))D_{t}P_{t}$ (B.25)
for some $(d\times d)$ diagonal matrix $D_{t}$ with ±$\pm 1$ diagonal elements and some $(d\times d)$ permutation matrix $P_{t}$ (that may vary in time). Rearranging the above equation gives
$\sum_{m=2}^{M}\alpha_{m,t}(B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})-B_{m}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m})D_{t}P_{t})=-\alpha_{1,t}B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})(I_{d}-D_{t}P_{t}).$ (B.26)
Suppose first $D_{t}P_{t}=I_{d}$. Equation (B.27) then simplifies to
$\sum_{m=2}^{M}\alpha_{m,t}(B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})-B_{m}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m}))=0.$ (B.27)
By Condition (B) of Proposition 2, the vector $(\alpha_{1,t},...,\alpha_{M,t})$ takes a value for some $t$ such that none of its entries is zero. At each such time period $t$, Equation (B.27) is nontrivial, as by invertibility, none of the columns of $B_{m}$ can be equal or negatives of each other. Hence, the set of matrices $B_{2},...,B_{M}$ that satisfy Equation (B.27) forms a subspace in $\mathbb{R}^{d\times d(M-1)}$ with the dimension of this subspace being less than $(M-1)d^{2}$, as it must satisfy a nontrivial linear equation. Consequently, the set of matrices $B_{2},...,B_{M}$ that satisfy Equation (B.27) has a Lebesgue measure zero in $[B_{2}:\cdots:B_{M}]\in\mathbb{R}^{d\times d(M-1)}$, implying that the set of matrices $B_{1},...,B_{M}$ satisfying this equation have a Lebesgue measure zero in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times Md}$.
Suppose then $D_{t}P_{t}\neq I_{d}$ and denote the left side of Equation (B.26) as $C_{t}\in\mathbb{R}^{d\times d}$. By Condition (B) of Proposition 2, the vector $(\alpha_{1,t},...,\alpha_{M,t})$ takes a value for some $t$ such that none of its entries is zero. At each such time period $t$, the right side of Equation (B.26) is nonzero, thus, implying that also $C_{t}$ is not a matrix zeros. Equation (B.26) is thereby nontrivial and can be expressed as
$B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})(I_{d}-D_{t}P_{t})=\alpha_{1,t}^{-1}C_{t}$ (B.28)
The set of matrices $B_{1}$ that satisfy Equation (B.28) forms a subspace in $\mathbb{R}^{d\times d}$ with the dimension of this subspace being less than $d^{2}$, as it must satisfy a nontrivial linear equation. It follows that the set of matrices $B_{1}$ that satisfy Equation (B.28) has Lebesgue measure zero on $\mathbb{R}^{d\times d}$, and therefore, the set of matrices $B_{1},...,B_{M}$ that satisfy Equation (B.28) has a Lebesgue measure zero in $[B_{1}:\cdots B_{M}]\in\mathbb{R}^{d\times Md}$.
By the above discussion, for any single transformation $D_{t}P_{t}$ of $B_{y,t}$, the set of matrices $B_{1},...,B_{M}$ that satisfy Equation (B.26) for any time period when all the entries of $(\alpha_{1,t},...,\alpha_{M,t})$ are nonzero has a Lebesgue measure zero in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times Md}$. Since there exists a finite number of such transformations and a countable union of measure zero sets has a measure zero, it follows that the set of matrices $B_{1},...,B_{M}$ that satisfy Equation (B.27) for all $t$ has a Lebesgue measure zero in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times Md}$. This concludes that, almost everywhere in $[B_{1},...,B_{M}]\in\mathbb{R}^{d\times Md}$, the matrix $B_{y,t}(\boldsymbol{\tau},\boldsymbol{\iota})$ cannot be obtained by reordering or changing the signs of the columns of $B_{y,t}(\tilde{\boldsymbol{\tau}},\tilde{\boldsymbol{\iota}})$ whenever $\boldsymbol{\tau}_{m}\neq\tilde{\boldsymbol{\tau}}_{m}$ or $\boldsymbol{\iota}_{m}\neq\tilde{\boldsymbol{\iota}}_{m}$ for any $m=2,...,M$
Next, we make use of the above result together with Lemma 2 to show that $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times Md}$. Consider the impact matrix $B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ with some invertible $(d\times d)$ matrices $B_{1},...,B_{M}$. By Lemma 2, $B_{y,t}$ is uniquely identified at each $t$ up to ordering and signs of its columns (conditionally on $\mathcal{F}_{t-1}$). Since the ordering and signs of the columns of $B_{1}$ are fixed (Condition (A) of Proposition 2), we then obtain from the above results that almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, reordering or changing the signs of the columns of $B_{2},...,B_{M}$ would, at some $t$, lead to an impact matrix that cannot obtained by reordering or changing the signs of the columns of $B_{y,t}$. Hence, models before and after reordering or changing the signs of the columns of any of $B_{2},...,B_{M}$ are not be observationally equivalent. The ordering and signs of the columns of $B_{1},...,B_{M}$ are thereby effectively fixed under Condition (A) of Proposition 2, implying that the ordering and signs of the columns of $B_{y,t}$ are also fixed for all $t$ (almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$). Thus, $B_{y,t}$ is uniquely identified at all $t$.
Since $B_{y,t}$ is uniquely identified and two models are observationally equivalent if and only if they are observationally equivalent at all $t$, a unique identification of the matrices $B_{1},...,B_{M}$ is achieved if an impact matrix that is identical to $B_{y,t}$ for all $t$ cannot be obtained with any other invertible $(d\times d)$ matrices $\tilde{B}_{1},...,\tilde{B}_{M}$. Suppose an impact matrix identical to $B_{y,t}$ for all $t$ is obtained with some invertible $(d\times d)$ matrices $\tilde{B}_{1},...,\tilde{B}_{M}$. Then,
$\sum_{m=1}^{M}\alpha_{m,t}(B_{m}-\tilde{B}_{m})=0,$ (B.29)
for all $t$.
Denote the $ij$th element of $\tilde{B}_{m}$ as $\tilde{b}_{ij}^{(m)}$. At each $t$, Equation (B.29) then defines a system of linear equations involving $d^{2}$ equations and $Md^{2}$ variables $\tilde{b}_{ij}^{(m)}$, $j,i=1,...,d$, $m=1,...,M$. The weights $\alpha_{m,t}$ are treated as constant coefficients here (some of which may be zero), as they are predetermined at each $t$ (by $\mathcal{F}_{t-1}$-measurability or nonrandomness), and the parameters in $B_{1},...,B_{M}$ are also constants here.
Each scalar equation in (B.29) is a linear equation in $M$ variables. Thus, each scalar equation in (B.29) defined together for all $t=1,...,T$ defines a system of $T$ linear equations and $M$ variables. By Condition (3) of Lemma 2 (assumed in Proposition 2), there exists $M$ indices $t_{1},...,t_{M}\in\{1,....,T\}$ such that the corresponding coefficient vectors $(\alpha_{1,t_{i}},...,\alpha_{M,t_{i}})$, $i=1,...,M$, are linearly independent. Hence, it follows from Lemma B.2 that there exists at most one solution to each such system of scalar equations, and thereby also to the system of $Td^{2}$ linear equations (B.29), $t=1,....,T$, in $Md^{2}$ variables. Since one solution is $\tilde{B}_{m}=B_{m}$ for all $m=1,...,M$, this is the unique solution, concluding that the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times dM}$, given that the transition weights $\alpha_{m,t}$, $m=1,...,M$, $t=1,...,T$, are uniquely identified. For exogenous models, the identification of the transition weights is given. For logistic models, by Lemma 2, the transition weights weights are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ and, if $z_{t}$ is exogenous, $(z_{1},...,z_{T})\in\mathbb{R}^{T}$, concluding the proof.∎
B.5 Proof of Theorem A.1
Saikkonen (2008) establishes stationarity, ergodicity, and mixing properties of a STVEC model by making use of the theory of Markov chains. Kheifets and Saikkonen (2020), in turn, show that the results of Saikkonen (2008) are applicable to their STVAR model, as it can be obtained as a special case of the STVEC model of Saikkonen (2008). However, the results of Saikkonen (2008) and Kheifets and Saikkonen (2020) are not directly applicable to our structural STVAR model, because our model defines the conditional covariance matrix differently to the STVEC model of Saikkonen (2008) and STVAR model of Kheifets and Saikkonen (2020). Hence, certain steps need to be taken show the applicability of the results of Saikkonen (2008) to our structural STVAR model.
Note that by Assumption A.1(II), Assumption 1 in Saikkonen (2008) is satisfied. Moreover, by the constraint $\sum_{m=1}^{M}\alpha_{m,t}=1$, Assumption 2(a) in Saikkonen (2008) satisfied, and as we assume that the set $\text{l}_{2}$ in Saikkonen (2008) is an empty set, also Assumption 2(b) is satisfied. Consequently, Condition (10) in Saikkonen (2008) is satisfied and by our Assumption A.1(I) also Condition (19). The difference of our setup to the conditions of Theorem 1 of Saikkonen (2008) is that our STVAR model specifies the conditional covariance matrix differently to the STVEC model defined in Equation (8) of Saikkonen (2008).
Besides the conditional covariance matrix, our structural STVAR model defined in Equations (3.2) and (3.3) is obtained from the stochastic process $z_{t}$ defined in Equation (17) (or equally (15)) of Saikkonen (2008) as follows ($z_{t}$ of Saikkonen (2008) should not be confused with the switching variable of this paper). First, we assume that the number of unit roots in the model is zero, so $n=r$ in Saikkonen (2008). Consequently, $z_{t}=y_{t}$ and, in the definition of the matrix $J$ defined above Equation (17) of Saikkonen (2008), $\beta=c=I_{d}$ (see Saikkonen, 2008, Note 2). The matrix $J$ is thereby nonsingular, and following Kheifets and Saikkonen (2020), we can choose the matrix $S$ in Equation (17) to be the inverse of $J^{\prime}$. Second, we assume the transition functions $h_{s}(S^{\prime}JZ_{t-1},\eta_{t})$ of Saikkonen (2008) are independent of the random variable $\eta_{t}$. Hence, the transition functions are functions of $Z_{t-1}=(y_{t-1},...,y_{t-p})$, and we can write $h_{s}(S^{\prime}JZ_{t-1},\eta_{t})=\alpha_{m,t}$. Using the notation of this paper, the first term on the right side of Equation (17) in Saikkonen (2008) then reduces to $\sum_{m=1}^{M}\alpha_{m,t}\sum_{i=1}^{p}A_{m,i}y_{t-i}$. Third, assume the set $\text{l}_{2}$ is empty, which implies that the second term on the right side of Equation (17), defined in Equation (11) of Saikkonen (2008), reduces to $\sum_{m=1}^{M}\alpha_{m,t}\phi_{m}$ (with the notation of this paper).
That is, the first two terms on the right side of Equation (17) in Saikkonen (2008), reduce to the first term on the right hand side of Equation (3.2) defining our structural STVAR model. The third term on the right side of Equation (17) in Saikkonen (2008), however, reduces to $(\sum_{m=1}^{M}\alpha_{m,t}\Omega_{m})^{1/2}e_{t}$, where $\Omega_{1},...,\Omega_{M}$ are positive definite $(d\times d)$ covariance matrices of the regimes. This is different to the second term of Equation (3.2), which takes the form $\sum_{m=1}^{M}\alpha_{m,t}B_{m}e_{t}$ when the impact matrix (3.3) is substituted to it. Nonetheless, it turns out that with only slight modifications, the proof of Theorem 1 in Saikkonen (2008) applies to our model as well.
The conditional covariance matrix is relevant in only a few places in the proof of Theorem 1 in Saikkonen (2008). Firstly, on page 312, Saikkonen (2008) makes use of the positive definiteness of the conditional covariance matrix $H(S^{\prime}JZ_{t-1})$ (or $\sum_{m=1}^{M}\alpha_{m,t}\Omega_{m}$ in our notation) to conclude that the smallest eigenvalue of $H(S^{\prime}Jx)$, $x\in\mathbb{R}^{dp}$, is bounded away from zero in compact subsets of $\mathbb{R}^{dp}$. Saikkonen (2008) argues that thereby by Theorem 2.2(ii) of Cline and Phu (1998), the Markov chain $Z_{t}$ is an irreducible and aperiodic T-chain on $\mathbb{R}^{dp}$, which is then used to conclude further results. Concerning our model, note that by Lemma 1, $B_{y,t}$ is invertible for all $t$ almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, implying that the conditional covariance matrix $B_{y,t}(\boldsymbol{y}_{t-1})B_{y,t}(\boldsymbol{y}_{t-1})^{\prime}$ is positive definite, where the dependence of the conditional covariance matrix on the $(dp\times 1)$ random vector $\boldsymbol{y}_{t-1}=(y_{t-1},...,y_{t-p})$ is made explicit in the notation. Hence, the smallest eigenvalue of $B_{y,t}(x)B_{y,t}(x)^{\prime}$, $x\in\mathbb{R}^{dp}$, is bounded away from zero in compact subsets of $\mathbb{R}^{dp}$. It follows from this, Equation (3.2), and the invertibility of $B_{y,t}$ that we can apply Theorem 2.2(ii) of Cline and Phu (1998) to conclude that the Markov chain $\boldsymbol{y}_{t}$ is an irreducible and aperiodic T-chain on $\mathbb{R}^{dp}$, similarly to the process $Z_{t-1}$ in Saikkonen (2008).
Secondly, the conditional covariance matrix appears in the proof of Saikkonen (2008) on page 313, where Saikkonen (2008) shows that an appropriate version of Condition (15.3) of Meyn and Tweedie (1993, p. 358) is satisfied. To show that this condition holds for our model, we need to show that an appropriate version of Equation (A.5) in Saikkonen (2008) holds for our model, i.e., that the term $E[\sup_{A\in\mathcal{A}^{j}}||AB_{y,t}(x)e_{t}||^{2}]$ is bounded from above by a finite real number, where $j$ is a positive integer, $\mathcal{A}=\{\boldsymbol{A}_{1},...,\boldsymbol{A}_{M}\}$, and $\boldsymbol{A}_{1},...,\boldsymbol{A}_{M}$ as well as the set $\mathcal{A}^{j}$ are defined in Appendix A.
Similarly to Saikkonen (2008, p. 313), clearly,
$\sup\{||A||:A\in\cup_{j=1}^{N}\mathcal{A}^{j}\}<c,$ (B.30)
where |||||$\cdot$||| | ⋅ | | is Eucledian norm, $N$ is a positive integer, and $c$ a finite real number. Hence, because $B_{y,t}(x)$ is a constant here, by Jensen’s inequality and submultiplicativity of norms, we have
$E\left[\sup_{A\in\mathcal{A}^{j}}||AB_{y,t}(x)e_{t}||^{2}\right]\leq\sup_{A\in\mathcal{A}^{j}}||A||^{2}||B_{y,t}(x)||^{2}E[||e_{t}||]^{2},$ (B.31)
where the right side is bounded from above by some finite constant (by (B.30) and the finite second moment of $e_{t}$ given by Assumption 1). Thus, an appropriate version of Equation (A.5) in Saikkonen (2008) holds for our model. Since the rest of the proof of Theorem 1 in Saikkonen (2008) thereby applies to our model as well, this concludes the proof.∎
Appendix C Penalized nonlinear least squares estimation of STVAR models
This appendix describes how the parameters of the structural STVAR model discussed in Section 3 can be estimated by the method of nonlinear least squares (NLS) as explained in Hubrich and Teräsvirta (2013, Section 4.2) for logistic STVAR models. After describing the NLS estimation, penalized nonlinear least squares (PNLS) estimation is considered by introducing a penalization term to the sum of squares of residuals function.
C.1 Nonlinear least squares estimation
Consider the following sum of squares of residuals function
$Q(\alpha)=\sum_{t=1}^{T}u_{t}(\alpha)^{\prime}u_{t}(\alpha),$ (C.1)
where $u_{t}(\alpha)=y_{t}-\sum_{m=1}^{M}\alpha_{m,t}\mu_{m,t}$ and the parameter $\alpha$ contains the weight function parameters. [^5] With threshold weights (2.3), $\alpha=(r_{1},...,r_{M-1})$, and with logistic weights (2.4) weights, $\alpha=(c,\gamma)$. If the transition weights are exogenous, we simply treat them as known constants. The NLS estimates of the parameters $\phi_{m},A_{m,j}$, $j=1,...,p$, corresponding to the weight function parameter $\alpha$ are obtained by minimizing $Q(\alpha)$ with respect to these parameters.
If the transition weights $\alpha_{1,t},...,\alpha_{M,t}$ are not exogenous, also the parameter $\alpha$ (they depend on) needs to be estimated. Since the NLS estimates of the AR parameters are analytically obtained for given $\alpha$, the optimization can be performed with respect to $\alpha$ only. However, this optimization problem is multimodal, and thereby standard gradient based optimization methods that are not able escape from local maxima are not sufficient. Since we only need approximate estimates to serve as starting values for Step 2 and 3 estimation in the three-step procedure described in Appendix 4.3, the following procedure is adopted.
First, determine the values of the transition weights $\alpha_{m,t}$, $m=1,...,M$, $t=1,...,T$, that can be deemed reasonable for estimation, i.e., transition weights that allocate a sufficient enough contribution to each regime in the model. As an extreme example, transition weights with $\alpha_{1,t}=0$ for all $t=1,...,T$ are not reasonable, as Regime 1 would not be involved in the model at all. Thus, each regime should contribute to the model to an extend that is reasonable enough for estimation.
One way to assess the contribution of each regime is by summing over the transition weights, i.e., calculating the sums $\dot{\alpha}_{m}\equiv\sum_{t=1}^{T}\alpha_{m,t}$, $m=1,...,M$. In the special case of discrete transition weights such as the threshold weights (2.3), the sums $\dot{\alpha}_{m}$ express the numbers of observations from each regime. Therefore, one way to determine whether a regime contributes enough to the model is to check whether the sums $\dot{\alpha}_{m}$ are large enough compared to the number of parameters in that regime for all $m=1,...,M$. Denote the large enough sum as $\tilde{T}$. We assume that if there are $k$ parameters in a regime, and there should be at least three times as many observations from all the $d$ variables combined, implying $\tilde{T}=3kd^{-1}$, then transition weights that satisfy $\dot{\alpha}_{m}\geq\tilde{T}$ for all $m=1,...,M$ are deemed reasonable enough. [^6] We acknowledge that our assumption of at least three times as many observations from all $d$ variables combined as there are parameters in a regime is somewhat arbitrary, but it attempts to deter obvious cases of overfitting, while still allowing flexibility when the model is large compared to the number of observations.
Second, specify a range of reasonable parameter values for each scalar parameter in the $(s\times 1)$ vector $\alpha$, and specify a grid of parameter values within that range. For instance, for a logistic STVAR model, the location parameter should, at least, lie between the lowest and greatest value in the observed series of the switching variable, whereas the scale parameter can range from a small positive number some large positive number that implies almost discrete regime-switches. With $n$ grid points for each $s$ scalar parameters in $\alpha$, the number of different parameter vectors $\alpha$ with values in the grid points is $n^{s}$. Hence, we recommend selecting a sufficiently small number of grid points $n$, so that estimation can be performed in a reasonable computation time. Denote the $i$th vector of $\alpha$ containing parameter values in the grid points as $\alpha^{(i)}$, $i=1,...,n^{s}$. Then, obtain the subset of the grid points $\alpha^{(i)}$ that satisfy $\dot{\alpha}_{m}\geq\tilde{T}$ for all $m=1,...,M$. Denote this subset as $\mathcal{W}$.
Finally, calculate the NLS estimates of the AR parameters for all $\alpha^{(i)}\in\mathcal{W}$, and the estimates yielding the smallest sum of squares of residuals $Q(\alpha)$ are the approximate NLS estimates of the parameters $\phi_{1},....,\phi_{M}$, $\varphi_{1},...,\varphi_{M}$, and $\alpha$. Notably, the NLS estimation does not necessarily produce estimates that satisfy the usual stability condition for all or any of the regimes.
C.2 Penalized nonlinear least squares estimation
Penalized nonlinear least squares estimation of STVAR models works similarly to the NLS estimation, except that instead of the sum of squares function (C.1), the penalized sum of squares function is minimized. We define the penalized sum of squares function as
$PQ(\alpha)=Q(\alpha)-P(\alpha),$ (C.2)
where the penalization term $P(\alpha)$ increases as the parameter values approach the boundary of, or move further into, an uninteresting region of the parameter space. To focus on avoiding estimates that do not satisfy the usual stability condition for all regimes, we define the penalty term as
$P(\alpha)=\kappa\widehat{RSS}\sum_{m=1}^{M}\sum_{i=1}^{dp}\text{max}\{0,|\rho(\boldsymbol{A}_{m}(\boldsymbol{\theta}))_{i}|-(1-\eta)\}^{2},$ (C.3)
where
$\widehat{RSS}=\underset{\alpha\in\mathcal{W}}{\min}Q(\alpha)$ (C.4)
is the minimized residual sum of squares, $|\rho(\boldsymbol{A}_{m}(\boldsymbol{\theta}))_{i}|$ is the modulus of the $i$th eigenvalue of the companion form AR matrix of Regime $m$, the tuning parameter $\eta\in(0,1)$ determines how close to the boundary of the stability region the penalization starts, and $\kappa>0$ determines the strength of the penalization. Whenever the companion form AR matrix of a regime has eigenvalues greater than $1-\eta$ in modulus, the penalization term (4.4) is greater than zero, and it increases in the modulus of these eigenvalues. The penalty term (C.3) is multiplied by $\widehat{RSS}$ to standardize the strength of the penalization with respect to the fit of the model.
The penalized NLS estimates of $\phi_{m},A_{m,j}$, $j=1,...,p$, and $\alpha$, are obtained by first minimizing $Q(\alpha)$ for all $\alpha\in\mathcal{W}$, then calculating the penalty term $P(\alpha)$ for all $(\alpha)\in\mathcal{W}$, and finally selecting the estimates that give the smallest minimized the penalized sum of squares (C.2). In our empirical analysis in Section 5, we use the tuning parameter values $\eta=0.05$ and $\kappa=0.2$, in line with the tuning parameter values used in the PML estimation (see Section 4.2).
Appendix D A Monte Carlo study
This appendix present a Monte Carlo study that assesses the performance of the proposed penalized maximum likelihood estimator and the three-step estimation procedure discussed in Section 4. We consider simple bivariate (structural) TVAR and LSTVAR models with $p=1$ and $M=2$, two specifications of each, with the first lag of the first variable as the switching variable. To maintain some constancy, we use the same impact matrices across all specifications, $\text{vec}(B_{1})=(0.6,-0.3,0.2,0.4)$ and $\text{vec}(B_{2})=(0.7,0.1,0.3,0.8)$, as well as the same degrees-of-freedom and skewness parameter values, $\nu_{1}=2.5$, $\nu_{2}=12.0$, $\lambda_{1}=-0.5$, and $\lambda_{2}=0.2$. For TVAR models, we use the threshold parameter value $r_{1}=0.8$ and for LSTVAR models the location and scale parameter values $c=0.8$ and $\gamma=5.0$.
For both the TVAR model and the LSTVAR model, we consider two specifications of the AR matrices, presented in Table D.1. In the first specification (labelled as TVAR 1 and LSTVAR 1 for the TVAR and LSTVAR models, respectively), the AR matrices are well inside the stability region (see Section 4.1). In the second specification (labelled as TVAR 2 and LSTVAR 2 for the TVAR and LSTVAR models, respectively), the AR matrices are well inside the region where penalization term is strictly positive, allowing us to assess the performance of our estimation method when the true parameter value lies in the penalization region. To illustrate the closeness to the boundary, the modulus of the roots of the companion form AR matrix are presented in Table D.1 (as in our empirical application, we use the tuning parameter values $\eta=0.05$ and $\kappa=0.2$ for the penalization term, so the penalization term is strictly positive when at least one the roots of the companion form AR matrix is larger than $0.95$ in modulus). The intercept parameter values are adjusted so that the unconditional means of the regimes are approximately equal in both of the specifications (see Table D.1), so that the main difference between the specifications is that the second one includes true parameter values from inside the penalization region.
Regime 1 Regime 2
$\phi_{1}$ $A_{1,1}$ $\mu_{1}$ $||\boldsymbol{A}_{1}||$ $\phi_{2}$ $A_{2,1}$ $\mu_{2}$ $||\boldsymbol{A}_{2}||$
[…]
Table D.1: The parameters values of $\phi_{1},\phi_{2},A_{1,1}$, and $A_{2,1}$ used for the models TVAR 1, TVAR 2, LSTVAR 2, and LSTVAR 2 in the Monte Carlo study. Also, the corresponding unconditional means of the regimes $\mu_{m}=(I_{d}-A_{m,1})^{-1}\phi_{m}$, $m=1,2$, and modulus of the eigenvalues of the companion form AR matrices of the regimes $||\boldsymbol{A}_{m}||$, $m=1,2$.
For each of the four model specifications, we generate $500$ samples of length $250,500,1000,2000$, and $10000$. Then, we fit the first-order two-regime TVAR model (for specifications TVAR 1 and 2) or LSTVAR model (for specifications LSTVAR 1 and 2) to each of the generated sample using the three-step estimation procedure based on the penalized log-likelihood function, proposed in Section 4. After estimation, the ordering and signs of the columns of $B_{1}$ and $B_{2}$ are standardized by assuming, without loss of generality, that the degrees-of-freedom parameters are in an ascending order and that the first skewness parameter is positive and the second one negative, in line with the true parameter values. To estimate the bias of the PML estimator, we calculate the average estimation errors $\hat{\boldsymbol{\theta}}-\boldsymbol{\theta}$ over the $500$ Monte Carlo repetitions, and to estimate the spread of the estimates, we calculate their sample standard deviations.
$T=250$ $T=500$ $T=1000$ $T=2000$ $T=10000$
TVAR 1 $\phi_{1}$ $0.01$ $(0.07)$ $0.01$ $(0.05)$ $0.00$ $(0.04)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
$0.00$ $(0.07)$ $0.00$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
$\phi_{2}$ $0.00$ $(0.16)$ $-0.01$ $(0.10)$ $0.00$ $(0.07)$ $0.00$ $(0.05)$ $0.00$ $(0.02)$
$0.01$ $(0.28)$ $0.00$ $(0.19)$ $0.00$ $(0.13)$ $0.00$ $(0.09)$ $0.00$ $(0.04)$
$A_{1,1}$ $-0.01$ $(0.08)$ $0.00$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
$-0.01$ $(0.10)$ $0.00$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
$0.00$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$
$-0.01$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
$A_{2,1}$ $-0.01$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
$-0.01$ $(0.11)$ $0.00$ $(0.07)$ $0.00$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$
$0.00$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$
$-0.01$ $(0.07)$ $-0.01$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
$B_{1}$ $0.20$ $(0.52)$ $0.09$ $(0.33)$ $0.05$ $(0.21)$ $0.02$ $(0.12)$ $0.01$ $(0.04)$
$-0.08$ $(0.34)$ $-0.04$ $(0.19)$ $-0.02$ $(0.11)$ $-0.01$ $(0.06)$ $0.00$ $(0.02)$
$-0.02$ $(0.09)$ $-0.01$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
$-0.03$ $(0.10)$ $-0.01$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
$B_{2}$ $0.30$ $(0.61)$ $0.14$ $(0.39)$ $0.07$ $(0.24)$ $0.03$ $(0.13)$ $0.00$ $(0.04)$
$0.04$ $(0.15)$ $0.02$ $(0.08)$ $0.01$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
$-0.01$ $(0.07)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$ $0.00$ $(0.00)$
$-0.03$ $(0.18)$ $-0.01$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
$r_{1}$ $0.00$ $(0.07)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.00)$ $0.00$ $(0.00)$
$\nu_{1}$ $0.05$ $(0.53)$ $0.02$ $(0.34)$ $0.01$ $(0.23)$ $0.00$ $(0.16)$ $0.00$ $(0.07)$
$\nu_{2}$ $165.03$ $(700.13)$ $105.25$ $(757.09)$ $13.98$ $(174.19)$ $1.48$ $(6.32)$ $0.16$ $(1.40)$
$\lambda_{1}$ $-0.03$ $(0.09)$ $-0.01$ $(0.06)$ $-0.01$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
$\lambda_{2}$ $0.01$ $(0.10)$ $0.00$ $(0.07)$ $0.00$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
TVAR 2 $\phi_{1}$ $-0.02$ $(0.12)$ $-0.01$ $(0.09)$ $-0.01$ $(0.06)$ $-0.01$ $(0.05)$ $0.00$ $(0.02)$
$-0.05$ $(0.14)$ $-0.04$ $(0.11)$ $-0.02$ $(0.07)$ $-0.02$ $(0.04)$ $-0.01$ $(0.02)$
$\phi_{2}$ $-0.02$ $(0.16)$ $-0.01$ $(0.10)$ $0.00$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.02)$
$-0.01$ $(0.30)$ $0.00$ $(0.18)$ $0.01$ $(0.12)$ $0.01$ $(0.08)$ $0.00$ $(0.04)$
$A_{1,1}$ $-0.02$ $(0.07)$ $-0.01$ $(0.04)$ $-0.01$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.00)$
$-0.01$ $(0.07)$ $-0.01$ $(0.05)$ $-0.01$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
$0.00$ $(0.07)$ $0.00$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
$-0.03$ $(0.06)$ $-0.02$ $(0.04)$ $-0.01$ $(0.03)$ $-0.01$ $(0.02)$ $-0.01$ $(0.01)$
$A_{2,1}$ $0.00$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$
$0.00$ $(0.10)$ $0.00$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
$-0.01$ $(0.04)$ $0.00$ $(0.02)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$
$-0.02$ $(0.07)$ $-0.01$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
$B_{1}$ $0.38$ $(0.64)$ $0.16$ $(0.40)$ $0.08$ $(0.24)$ $0.04$ $(0.12)$ $0.01$ $(0.04)$
$-0.15$ $(0.43)$ $-0.07$ $(0.25)$ $-0.04$ $(0.15)$ $-0.02$ $(0.08)$ $0.00$ $(0.03)$
$-0.04$ $(0.12)$ $-0.02$ $(0.08)$ $-0.01$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
$-0.05$ $(0.18)$ $-0.02$ $(0.07)$ $-0.01$ $(0.04)$ $-0.01$ $(0.03)$ $0.00$ $(0.01)$
$B_{2}$ $0.55$ $(0.73)$ $0.24$ $(0.44)$ $0.12$ $(0.25)$ $0.05$ $(0.12)$ $0.01$ $(0.04)$
$0.06$ $(0.17)$ $0.03$ $(0.09)$ $0.01$ $(0.05)$ $0.01$ $(0.03)$ $0.00$ $(0.01)$
$-0.01$ $(0.13)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$ $0.00$ $(0.00)$
$-0.05$ $(0.25)$ $-0.01$ $(0.08)$ $0.00$ $(0.02)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
$r_{1}$ $0.08$ $(0.45)$ $0.02$ $(0.24)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.00)$
$\nu_{1}$ $0.05$ $(2.77)$ $-0.07$ $(0.29)$ $-0.05$ $(0.21)$ $-0.03$ $(0.15)$ $0.00$ $(0.07)$
$\nu_{2}$ $107.17$ $(542.02)$ $93.32$ $(1273.27)$ $10.09$ $(81.02)$ $1.29$ $(5.10)$ $0.09$ $(1.47)$
$\lambda_{1}$ $-0.04$ $(0.11)$ $-0.02$ $(0.05)$ $-0.01$ $(0.04)$ $-0.01$ $(0.03)$ $0.00$ $(0.01)$
$\lambda_{2}$ $-0.01$ $(0.12)$ $0.00$ $(0.07)$ $0.00$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
Table D.2: Results for the structural TVAR models from the Monte Carlo study assessing the performance of our penalized log-likelihood based three-step estimation method based on the PML estimator discussed in Section 4. The average estimation error is reported for the samples of length $250$, $500$, $1000$, $2000$, and $10000$ in each column first, and next to the biases are the standard deviations of the estimates in parentheses. The results are based on $500$ Monte Carlo repetitions.
$T=250$ $T=500$ $T=1000$ $T=2000$ $T=10000$
LSTVAR 1 $\phi_{1}$ $0.01$ $(0.17)$ $0.01$ $(0.08)$ $0.01$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.02)$
$0.00$ $(0.25)$ $0.00$ $(0.12)$ $0.00$ $(0.07)$ $0.00$ $(0.05)$ $0.00$ $(0.02)$
$\phi_{2}$ $-0.01$ $(0.17)$ $-0.01$ $(0.11)$ $0.00$ $(0.08)$ $0.00$ $(0.06)$ $0.00$ $(0.03)$
$-0.02$ $(0.39)$ $-0.02$ $(0.24)$ $-0.01$ $(0.15)$ $0.00$ $(0.11)$ $0.00$ $(0.05)$
$A_{1,1}$ $0.00$ $(0.15)$ $0.01$ $(0.07)$ $0.00$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
$-0.02$ $(0.26)$ $-0.01$ $(0.13)$ $0.00$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.01)$
$0.02$ $(0.08)$ $0.01$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
$-0.01$ $(0.10)$ $-0.01$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
$A_{2,1}$ $0.00$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
$0.00$ $(0.14)$ $0.00$ $(0.09)$ $0.00$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.02)$
$-0.01$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$
$-0.01$ $(0.08)$ $-0.01$ $(0.05)$ $0.00$ $(0.04)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
$B_{1}$ $0.24$ $(0.60)$ $0.10$ $(0.37)$ $0.04$ $(0.20)$ $0.02$ $(0.12)$ $0.01$ $(0.04)$
$-0.09$ $(0.46)$ $-0.05$ $(0.24)$ $-0.02$ $(0.13)$ $-0.01$ $(0.08)$ $0.00$ $(0.03)$
$-0.03$ $(0.12)$ $-0.02$ $(0.08)$ $-0.01$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
$-0.05$ $(0.14)$ $-0.03$ $(0.08)$ $-0.01$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
$B_{2}$ $0.44$ $(0.72)$ $0.19$ $(0.42)$ $0.08$ $(0.22)$ $0.04$ $(0.12)$ $0.01$ $(0.04)$
$0.05$ $(0.17)$ $0.02$ $(0.09)$ $0.01$ $(0.05)$ $0.01$ $(0.03)$ $0.00$ $(0.01)$
$-0.01$ $(0.07)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$ $0.00$ $(0.00)$
$-0.03$ $(0.19)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
$c$ $0.03$ $(0.19)$ $0.00$ $(0.07)$ $0.00$ $(0.04)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
$\gamma$ $28.86$ $(121.66)$ $2.90$ $(25.52)$ $0.09$ $(0.75)$ $0.05$ $(0.47)$ $0.01$ $(0.19)$
$\nu_{1}$ $-0.02$ $(0.49)$ $-0.02$ $(0.32)$ $-0.02$ $(0.21)$ $-0.01$ $(0.15)$ $0.00$ $(0.07)$
$\nu_{2}$ $152.80$ $(597.75)$ $43.27$ $(266.76)$ $8.00$ $(52.02)$ $1.46$ $(5.55)$ $0.11$ $(1.37)$
$\lambda_{1}$ $-0.04$ $(0.09)$ $-0.02$ $(0.05)$ $-0.01$ $(0.04)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
$\lambda_{2}$ $0.02$ $(0.10)$ $0.01$ $(0.07)$ $0.00$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
LSTVAR 2 $\phi_{1}$ $-0.01$ $(0.18)$ $-0.01$ $(0.11)$ $-0.01$ $(0.08)$ $-0.01$ $(0.05)$ $0.00$ $(0.02)$
$-0.10$ $(0.19)$ $-0.06$ $(0.13)$ $-0.03$ $(0.08)$ $-0.02$ $(0.06)$ $-0.02$ $(0.02)$
$\phi_{2}$ $0.07$ $(0.67)$ $0.01$ $(0.11)$ $0.00$ $(0.07)$ $0.00$ $(0.05)$ $0.00$ $(0.02)$
$0.12$ $(1.18)$ $0.02$ $(0.19)$ $0.01$ $(0.13)$ $0.01$ $(0.08)$ $0.01$ $(0.04)$
$A_{1,1}$ $-0.05$ $(0.13)$ $-0.02$ $(0.07)$ $-0.01$ $(0.03)$ $-0.01$ $(0.02)$ $0.00$ $(0.01)$
$-0.03$ $(0.11)$ $-0.02$ $(0.08)$ $-0.01$ $(0.04)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
$0.03$ $(0.13)$ $0.01$ $(0.08)$ $0.00$ $(0.04)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
$-0.03$ $(0.08)$ $-0.03$ $(0.06)$ $-0.02$ $(0.03)$ $-0.01$ $(0.02)$ $-0.01$ $(0.01)$
$A_{2,1}$ $-0.02$ $(0.11)$ $-0.01$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$
$-0.02$ $(0.21)$ $-0.01$ $(0.06)$ $-0.01$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
$-0.01$ $(0.06)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$
$-0.03$ $(0.13)$ $0.00$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
$B_{1}$ $0.44$ $(0.74)$ $0.16$ $(0.41)$ $0.08$ $(0.24)$ $0.03$ $(0.12)$ $0.01$ $(0.05)$
$-0.14$ $(0.55)$ $-0.06$ $(0.27)$ $-0.04$ $(0.18)$ $-0.02$ $(0.09)$ $0.00$ $(0.03)$
$-0.04$ $(0.13)$ $-0.02$ $(0.08)$ $-0.01$ $(0.06)$ $-0.01$ $(0.03)$ $0.00$ $(0.01)$
$-0.03$ $(0.19)$ $-0.02$ $(0.09)$ $-0.02$ $(0.05)$ $-0.01$ $(0.04)$ $0.00$ $(0.02)$
$B_{2}$ $0.64$ $(0.84)$ $0.25$ $(0.44)$ $0.12$ $(0.25)$ $0.05$ $(0.12)$ $0.01$ $(0.04)$
$0.08$ $(0.19)$ $0.03$ $(0.09)$ $0.01$ $(0.05)$ $0.01$ $(0.03)$ $0.00$ $(0.01)$
$-0.01$ $(0.09)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$ $0.00$ $(0.00)$
$-0.04$ $(0.23)$ $-0.01$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
$c$ $0.23$ $(0.80)$ $0.05$ $(0.31)$ $0.00$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.01)$
$\gamma$ $79.82$ $(349.54)$ $22.77$ $(90.88)$ $5.71$ $(44.05)$ $0.39$ $(4.28)$ $0.10$ $(0.30)$
$\nu_{1}$ $-0.11$ $(0.57)$ $-0.07$ $(0.30)$ $-0.05$ $(0.21)$ $-0.03$ $(0.15)$ $0.00$ $(0.07)$
$\nu_{2}$ $192.21$ $(889.55)$ $38.30$ $(233.50)$ $9.58$ $(79.39)$ $1.37$ $(4.71)$ $0.15$ $(1.42)$
$\lambda_{1}$ $-0.05$ $(0.09)$ $-0.02$ $(0.06)$ $-0.01$ $(0.04)$ $-0.01$ $(0.03)$ $0.00$ $(0.01)$
$\lambda_{2}$ $0.00$ $(0.11)$ $0.00$ $(0.07)$ $0.00$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
Table D.3: Results for the structural LSTVAR models from the Monte Carlo study assessing the performance of our penalized log-likelihood based three-step estimation method based on the PML estimator discussed in Section 4. The average estimation error is reported for the samples of length $250$, $500$, $1000$, $2000$, and $10000$ in each column first, and next to the biases are the standard deviations of the estimates in parentheses. The results are based on $500$ Monte Carlo repetitions.
The results of the Monte Carlo study are presented in Table D.2 for the TVAR models and in Table D.3 for the LSTVAR models. The results show that there is some but mostly very small bias, which diminishes when the sample size increases. Since also the standard deviation of the estimates diminishes with the sample size, our results are in line with consistency of the PML estimator. Nonetheless, for certain parameters, there appears to be a notable bias in small samples.
First, the estimator of the second degrees-of-freedom parameter (true parameter value $\nu_{2}=12$) is substantially biased upwards in samples of length $T=250,500$, and $1000$, and also the standard deviation of the estimates is very large. This can explained by the property of the skewed $t$-distribution that when the degrees-of-freedom parameter value is already large, increasing it more has virtually no effect on the shape of the distribution. Consequently, if a large degrees-of-freedom parameter fits the data, the numerical optimization algorithms are incapable of discriminating between large and very large parameter values. Hence, possibly very large degrees-of-freedom parameter values drawn by random in Step 2 estimation may remain large through Step 3 estimation. Since it is clear that there is virtually no difference in the economic interpretations of large (say, $100$) and very large (say, $1000$) degrees-of-freedom parameter values, the large upward bias should not be a concern.
Second, in the LSTVAR models, there is a substantial upward bias also for the scale parameter $\gamma$ when the sample size is small. A large scale parameter value indicates almost discrete regime-switches, suggesting that the PML estimator is, in small samples, biased towards estimating too fast regime-switches when they are smooth in the true data generating process ($\gamma=5$). The upward bias is possibly heightened by the property that when $\gamma$ is large, the regime-switches are already nearly discrete, making it difficult for numerical estimation algorithms to properly distinguish between large and very large values, particularly in small samples.
Third, the estimators of the first elements of the first rows of $B_{1}$ and of $B_{2}$ exhibit a notable upward bias when the sample is small. A possible explanation is that the bias is related to the heavy tails of the skewed $t$-distribution, as the true degrees-of-freedom parameter value is small ($\nu_{1}=2.5$), and also, the estimators of the first elements of the second rows of $B_{1}$ and $B_{2}$ are biased more than the estimators of most of the other parameters. In particular, the estimates may be sensitive to extreme values drawn for the heavy-tailed skewed $t$ shocks, potentially amplifying small-sample bias.
In general, the small-sample bias of the PML estimator seems to be slightly larger the specifications where the true AR matrices are in the penalization region (TVAR 2 and LSTVAR 2) than the specifications where the true AR matrices lie well inside the stability region (TVAR 1 and LSTVAR 1). The differences are, however, only small, and even in small samples, there does not appear to be any particular issues with the estimation accuracy of the AR matrices lying inside the penalization region. Also, the TVAR models seem to be slightly more accurately estimated than the LSTVAR models, possibly because, unlike in TVAR models, in LSTVAR models also the smoothness of regime-switches needs to be estimated.
Appendix E Details on the empirical application
The estimation of the models, residual diagnostics, and computation of the GIRFs are carried out with the accompanying R package sstvars (Virolainen, 2025a) in which the introduced methods have been implemented. The R package sstvars also contains the dataset studied in the empirical application. In the below residual based model diagnostics, we have not labelled the residual series by any of the variables, because labelling them effectively requires labelling the corresponding structural shocks, and only the CPU shock has been labelled.
To study whether our model adequately captures the autocorrelation structure of the data, we depict the sample autocorrelation functions (ACF) and cross-correlation functions (CCF) of the (non-standardized) residuals for the first $24$ lags in Figure E.1. There is not much autocorrelation left in the residuals, but there is a particularly large correlation coefficient (CC) in the lag zero CCFs between the fourth and fifth residual series. In addition, there are a few moderately sized CCs in the ACFs and CCFs of some of the residual series, but the vast majority of the CCs are small. Hence, with the exception of the single large CC in the lag zero CCFs of the fourth and fifth residual series, our LSTVAR model appears to capture the autocorrelation structure of the data reasonably well.
To study our model’s adequacy to capture the conditional heteroscedasticity in the data, we depict the sample ACFs and CCFs of squared standardized residuals for the first $24$ lags in Figure E.2. There is a very large CC in the lag zero CCFs between the second and third residual series, which, however, disappears if the COVID-19 residual from March 2020 is excluded (not shown). In addition, there are several sizeable CCs in the ACFs and CCFs of some of the residual series. Thus, there are some inadequacies in capturing the conditional heteroscedasticity, but with the exception of the large CC in CCFs between the second and third residual series (due to the extreme COVID-19 volatility), the inadequacies do not seem particularly severe.
Finally, the quantile-quantile-plots based on the independent skewed $t$-distributions with mean zero, variance one, and degrees-of-freedom as well as skewness parameter values given by their estimates are presented in Figure E.3 for each residual series. They show that the marginal distributions of the series are reasonably well captured, although there is a single large outlier in the third residual series (which disappears if the March 2020 residual is removed; not shown). Therefore, the overall adequacy of the model seems, in our view, reasonable.
Figure E.1: Auto- and crosscorrelation functions of the residuals of the fitted two-regime sixth-order LSTVAR model for the lags $0,1,...,24$. The lag zero autocorrelation coefficients are omitted, as they are one by convention. The dashed lines are the $95\%$ bounds ±$\pm 1.96/\sqrt{T}$ for autocorrelations of IID observations.
Figure E.2: Auto- and crosscorrelation functions of the squared standardized residuals of the fitted two-regime sixth-order LSTVAR model for the lags $0,1,...,24$. The lag zero autocorrelation coefficients are omitted, as they are one by convention. The dashed lines are the $95\%$ bounds ±$\pm 1.96/\sqrt{T}$ for autocorrelations of IID observations.
Figure E.3: Standardized residual time series and quantile-quantile-plots of the fitted two-regime sixth-order LSTVAR model. The quantile-quantile plots are based on $t$–distributions with zero mean, variance one, and degrees-of-freedom as well as skewness parameter values given by their estimates.